### v1

# Can the 3 neutrino masses really be found using SN 1987A data?

Abstract
Neutrino masses remain a significant unsolved problem in physics and their nonzero value proves the Standard Model is incomplete. Currently, the values of the three masses only have upper limits from cosmology and experiments like KATRIN. This paper shows that the SN 1987A neutrino data can remrkably yield values for the three neutrino masses, and not merely upper limits. Although this seemingly preposterous idea was suggested a dozen years ago by the author, here it is demonstrated in a much more convincing manner with many new elements, including a stronger statistical treatment, a theoretical linkage to possible CPT violation, and most importantly, a thorough explanation of why the method used to find the three masses from supernova SN 1987A neutrino data really works. The key to finding the three neutrino masses is realizing why three normally accepted assumotions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any masses $1eV/c^{2}$ would be inconsistent with upper limits from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A is too great for the method to work.
Keywords: SN 1987A, Neutrino, supernova, CPT violation, tachyon, KATRIN
I Introduction
About $10^{58}$ neutrinos were emitted from the dying star Sanduleak 69 202 when it became the first supernova seen in 1987 (hence the name SN 1987A). Of those $10^{58}$ neutrinos only a couple of dozen were captured by Earth’s detectors. Yet, thousands of papers have been published about SN 1987A, and its well-studied nature is not surprising given that it is the only supernova in our galaxy since neutrino detectors first existed. The early articles on the neutrino mass obtained from SN 1987A data mostly, with one exception,(Cowsik1988, ) set upper limits on its value. (Arnett1987, ; Burrows1988, ) The upper limits restriction is based on the almost universal belief that the differences in the three neutrino mass states are simply too small to use the data to find the individual masses themselves.
Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found. As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) an extensive discussion of CPT violation, (2) a much improved statistical analysis, (3) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A.
I.1 Are $m^{2}<0,$ and $v>c$ Tachyons “unphysical”?
Many theorists have regarded quantum fields giving rise to physical tachyons as being either impossible or unlikely in light of the many theoretical difficulties they create. However, unobservable $m^{2}<0$ particles seem to be OK with theorists. Thus, for most theorists the idea that symmetry-breaking in the interaction of Higgs fields with massless gauge fields leads to the production of unobservable tachyons is not objectionable. The difficulties with “physical” tachyons having an $m^{2}<0$ that can be observed include their energy spectrum being unbounded from below, their frame-dependent and unstable vacuum state, and their noncovariant commutation rules. Nevertheless, there are also theorists who are accepting of physical tachyons, and recently Paczos et al. (Paczos2024, ) have shown that a covariant framework exists which allows for the proper quantization of tachyon fields which eliminates all such “unphysical” issues. Charles Schwartz is another theorist who has challenged all the reasons why tachyons are considered unphysical. (Schwartz2022, ).
I.2 The meaning of negative searches for tachyons
The question of whether $m^{2}<0$ and $v>c$ observable particles really exist is not theoretical, but strictly empirical. All reports of tachyons, including the 2011 OPERA experiment initial $v>c$ result, (Brumfiel2012, ) have later been shown to be incorrect. (Ehrlich2022, ) Most physicists therefore believe that the tachyon hypothesis can be dismissed, or at least relegated to the highly implausible category. A more reasonable conclusion is that as long as not a single neutrino has ever been measured to have either $m^{2}>0$ or $v<c,$ the tachyonic neutrino hypothesis remains viable. A decline in belief in physical tachyons is perfectly understandable as a psychological reaction to negative results or even worse, false positive results like OPERA. However, negative and false positive results do not suggest tachyons do not exist, only that the $m^{2}$ for neutrinos is very close to zero, so we cannot tell if $m^{2}>0$ or $m^{2}<0$. Neutrinos of course are the only known particles that might be tachyons since all others have been observed travelling at speeds $v<c,$ and tachyons if they exist must always be superluminal. Negative results for tachyonic neutrinos are not unlike negative results for searches for extraterestrials (SETI). In both cases, an appropriate response to a negative search would be “keep looking using more sensitive techniques.” Surprisinfly, however, if you ask an AI whether negative searches for tachyons make it more likely they don’t exist and then ask the same question about ET’s existence, you are illogically told yes for the former question and no for the latter.
I.3 The KATRIN experiment
The KATRIN experiment seeks to find the “effective mass” of the electron antineutrino $m^{2}=\Sigma|U_{i,j}|^{2}m_{j}^{2}.$ Its best value as of 2025 is negative: $m^{2}=-0.14^{+0.13}_{-0.15}eV^{2}.$ (KATRIN2024, ) This result, of course, leaves the true sign of $m^{2}$ ambiguous given the size of the uncertainty. Is it just a coincidence that the KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass value the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}?$ Ehrlich2015 Maybe it is a just a weird coincidence, but fortunately, as KATRIN takes more data the statistical uncertainty on the $m^{2}$ value now $0.108eV^{2}$ will be reduced. Whatever the final KATRIN results, recall that the experiment measures the effective mass of the electron neutrino and not the three individual masses comprising it, which is the primary interest of this paper.
II 3 Neutrino masses and SN 1987A
It is widely believed that we can set an upper limit on the neutrino masses based on oscillation experiments. These experiments yield values for separations in the mass squared of pairs of mass states: $\Delta m^{2}_{j,k}=m_{j}^{2}-m_{k}^{2},$ the two well-established ones being the atmospheric and solar $m^{2}$ splittings $\Delta m^{2}_{atm}$ and $\Delta m^{2}_{sol}.$
II.1 A true upper limit on Neutrino masses?
The three masses $m^{2}_{k}$ values are normally assumed to be no greater than $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}\approx 0.0025eV^{2}.$ It is important to recognize this upper limit is based on a questionable assumption, namely the non-existence of any third observed oscillation frequency. In other words, since the two oscillation frequencies found for (1) $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}$ and (2) $\Delta m^{2}_{atm}$ are so close, they are empirically indistinguishable. If a third observed oscillation frequency did exist (perhaps the result of a fourth sterile neutrino) then the assumption of the upper limit $m_{k}^{2}<0.0025eV^{2}$ would be false. In that case, it is conceivable the neutrino masses are large enough to be measurable based on the neutrino travel time over a very large distance. Moreover, the best place to measure the three masses based on travel time might be a galactic supernova, provided the neutrino emissions from it are nearly simultaneous. No other source of a neutrino burst, even that produced by an atomic bomb, could come close to the precision made possible by the enormous travel distance to a supernova.
II.2 Supernova SN 1987A neutrinos
Three neutrino detectors operated at the time supernova SN1987A occurred at 7:35 UT on Feb 23, 1987. Those detectors (Kamiokande II, IMB, and Baksan) recorded bursts of 12, 8 and 5 antineutrinos respectively at approximately the same time, and they recorded the visible energy, incoming direction and arrival times of the neutrinos. The three detectors were not perfectly synchronized, with a time difference between them being perhaps a few seconds. This lack of synch is partially removed in most analyses by assuming the first event in each detector occurred at time $t=0.$ The arriving electron antineutrinos were detected based on inverse beta decay
$\bar{\nu_{e}}+p\rightarrow n+e^{+}$ (1)
so that the antineutrino energy was actually 1.3 MeV higher than the recorded visible $(e^{+})$ energy. Those three neutrino bursts occurred about three hours before the star emitting them suddenly brightened. This late light arrival, of course does not imply that neutrinos are superlumnial. Rather, the three-hour gap between the neutrino burst and the later arrival of visible light, aligns with the small interaction between neutrinos with matter. Thus, photons but not neutrinos are delayed by interactions with the supenova ejecta.
II.3 The 5 hour early LSD neutrino burst
In addition to the three detectors, there was a fourth small detector known as LSD ln 1987 which was located under Mont Blanc. LSD reported a burst of five neutrinos which occurred 484 min before the other three detectors’ bursts. This early burst is therefore considered by most physicists to be a background fluctuation unrelated to SN 1987A because of its five hour early arrival. If the LSD burst were genuine it could constitute evidence that those neutrinos were tachyons, although other explanations also exist, including having SN1987A “bang” twice (Galeotti2016, ) and other ones ones as well. (Agafonova2018, ) Nevertheless, even raising the possibility of tachyons probably was enough to dismiss the existence of the LSD neutrino burst as unrelated to SN 1987A in the minds of most physicists.
The five neutrinos observed during a 7 second time interval in LSD would if spurious represent a very significant departure from background (roughly one neutrino every 80 seconds), and that background had been steady for months. In fact, the probability of a chance fluctuation giving rise to the LSD signal is extremely low, less than $1.4\times 10^{–6}.$ (Galeotti2016, ) Many observers also reject the LSD burst because no such early burst was seen in the three other detectors. However, the LSD neutrinos were all very low energy, and the other detectors could not have seen an LSD-like burst at that early time because of their higher energy thresholds. Thus, for example, the largest detector, Kamiokande II, had an energy threshold of 7 MeV, well above the energies of four of the five LSD neutrinos constituting the burst Another suspicious fact is that the main t=0 burst at 7:35 UT was not seen in LSD. However, that can also be explained because the small LSD detector did observe two neutrinos that time (not enough to qualify as a burst). Two final major problem with the LSD burst being associated with SN 1987A are (1) why the very low energies of all the 5 neutrinos comprising it ?, and (2 )why the virtual monochromaticity of the burst? In what follows we accept that the LSD neutrino burst was associated with SN 1987A, because all five above objections to it are answerable, and we address the monochromaticity problem later.
III Finding $m^{2}$ of SN 1987A neutrinos
The possibility of using the SN 1987A data to find the three neutrino masses may seem implausible, but we shall see that it can be done if the neutrinos were emitted from the star close to simultaneously. Let us first see why for exactly simultaneous neutrino emissions the neutrino arrival time relative to a photon directly determines their travel time (relative to light), and how using the measured neutrino energy $E,$ and arrival time t we can compute the mass of $\emph{i}ndividual$ neutrinos. In what follows $T=168,000$ years is the time photons required to reach us from SN 1987A.
If a photon’s speed en route to Earth is $c=D/T$ then a neutrino’s speed would be $v=D/(t+T)\approx c(1-t/T).$ Here t is the time delay of a neutrino relative to a photon. But since there can also be such a delay if a neutrino is emitted later than other neutrinos, so the definition of t only makes sense if almost all neutrinos were emitted during a very small time window. For the neutrino speed we have v/c=1t/T.v/c=1-t/T. and also $v/c=\sqrt{1-m^{2}c^{4}/E^{2}}\approx 1-m^{2}c^{4}/2E^{2}.$ Setting the two versions of $v/c$ equal
$v/c=1-t/T=1-m^{2}c^{4}/2E^{2}$ (2)
yields
$\frac{1}{E^{2}}=\big(\frac{2}{Tm^{2}c^{4}}\big)t\equiv Mt$ (3)
Eq. 3 implies that if a number of neutrinos with the same mass $m^{2}c^{4}$ were present and they had different arrival times t and energies E, then on a plot of $1/E^{2}$ versus $t,$ the neutrinos will be found to lie on or near a straight line through the origin of slope M given by
$M=\frac{2}{Tm^{2}c^{4}}$ (4)
Thus, if the SN 1987A data show such a clustering about a straight line, from the best-fit slope M we can find the common mass of all neutrinos on or near the line as:
$m^{2}c^{4}=\frac{2}{TM}$ (5)
Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line having negative slope.
To the extent that neutrinos all started out from SN 1987A at nearly the same instant, then since there are known to be three (and only three) neutrino masses, every neutrino should lie on or close to one of three straight lines passing through the origin. Any neutrinos that do not do so must either be a background event unrelated to SN 1987A, or else the result of non-simultaneous emissions. It is however possible that the data points will all be found to be associated with only one or two straight lines rather than three lines if two or three of the neutrino masses are too close to distinguish from one another.
Before we examine what the data show, there is a philosophical issue to mention. Given that in the standard view neutrino masses are less than $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}$ – a value so small that any spread in neutrino travel times could not possibly be large enough to observe the neutrino mass, we have two choices on how to proceed. The first is not to even bother to see what the data suggest (the ”Cremonini option”) given that we know any positive result must be an anomaly, given the small upper limt on the neutrino mass . However, we have made the second choice, namely not to make any assumptions on the size of the neutrino masses and be guided by what the data themselves show.
IV Fitting 30 data points to 3 lines
Strikingly, when the SN 1987A data is plotted (see Fig.1) every one of the 30 neutrinos lies on one of three straight lines through the origin, and very few are next to two lines. The least square fit of the 30 data points to three lines through the origin has two free parameters. The first parameter is $S$ which adjusts the trial slopes of the three straight lines in Fig. 1 simultaneously and the second free parameter H is the size of the constant horizontal error bar on each data point in Fig.1. Since each data point has both x and y error bars, we can then convert the x error into an ”effective” y error by multiplying the x uncertainty by the slope of the line, and combine it with the y uncertainty in quadrature. If a data point lies close to two straight lines, we simply choose the line that gives the smaller contribution to chi square. However, only a few of the data points near the origin are within $1\sigma$ of more than one straight line. The good fit shown in Fig 1 has a chi square of 25.4. Given 28 degrees of freedom we find $p=61\%.$ The excellent best fit was for a horizontal error bar of two seconds on all data points, and the fit would still be acceptable but significantly worse for a one second horizontal error bar ($p=18\%$.) For a fuller desccription of how the $\chi^{2}$ fit probability depends on the size of the horizontal error bar see Fig. 2 We have seen that exactly simultaneous neutrino emissions requires the data must lie on one of three lines through the origin in a $1/E^{2}$ versus t plot. Therefore the goodness of fits using 1 and 2 sec horizontal error bars shows they will continue to lie on or near these lines if the time window of neutrino emissions is no more than 1-2 seconds.
Figure 1: Best fit of 30 data points $(t,1/E^{2}),$ one for each neutrino to three straight lines that pass through the origin. Error bars in both horizontal and vertical are included in the fit. The best fit has $\chi^{2}=25.0$ for 28 dof, which yields $p=61\%$. The three best fit masses $m_{j}$ are obtained from the three best fit slopes according to Eq. 5. Note the different scales used for $t>0$ and $t<0$ The five LSD data points at t=17ksect=-17ksec have been slightly separated from each other for clarity.
The three masses corresponding to the slopes of the three straight lines are $m_{1}=22.4eV/c^{2}$ and $m_{2}=2.7eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ for the $m^{2}<0$ line. The big difference from the author’s previously published values of this fit is in the value of $m_{2}$ because of the past omission of horizontal error bars, which are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line. We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.7eV/c^{2}$ and the $2.7eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration, as well as its non-observation by KATRIN. (Serebrov2021, ; KATRIN2025, ). Of course, our $m_{1},m_{2},m_{3},$ however, are not sterile neutrino masses, but the three active neutrino masses.
Figure 2: The probability from the least squares fit to three straight lines through the origin versus the size of the horizontal error bars, assumed to be the same for all 30 data points. Obviously for $H>>5$ sec error bars the fit has $p\approx 100\%$
V Discussion
In this section we discuss eleven issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0,$ really does imply three specifc neutrino masses, and it is not just a mirage or background fluctuation.
V.1 Why not not exclude the LSD data?
Initially, the author rejected the LSD data. This rejection was not because it proviided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous. However since this and all the other reasons many people have for rejecting the LSD data can be answered, there really is no valid reason not to keep it.
V.2 Why are the LSD data bunched in time?
If the LSD burst of 5 neutrinos is genuine, it looks strange in Fig. 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec. The three straight line result would be more believable if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line. But recall that the $t<0$ data has a kilosecond time scale. If the five LSD data points were spread out along the negative sloped line and not bunched in time they never could have been distinguished from background because researchers were looking for a bunch of events in a seconds-long time interval.
V.3 Why do the LSD neutrinos have the same E?
The apparent near-constant energy of the LSD burst remains mysterious even to the physicists who accept its existence, but it is natural if the 5 neutrinos (and all others) must lie on one of three straight lines. Moreover, there is a theoretical basis for the monochromaticity of the LSD burst, as shown in several of the author’s publications that have provided independent evidence for it. That evidence was in the form of a proposed model for an 8 MeV neutrino line from a supernova, and further evidence related to the observed spectrum of MeV $\gamma$-rays from the galactic center. (Ehrlich2021, )
V.4 Alternate explanations for the LSD neutrinos?
All the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems. Thus,
  • If it was a statistical fluctuation it was a really big one $p=1.4\times 10^{-6}$
  • If it was due to a double bang the pattern in Fig. 1 should look very different as discussed in section 6.
  • If it was a genuine, low-energy pulse (e.g., from neutron capture on iron slabs around the detector), it would require the supernova to release an unusually high amount of energy in this form.
  • If it was due to 8 MeV dark matter as suggested in ref. (Ehrlich2021, ) this would still creates the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD neutrinos falling on the $m^{2}<0$ line.
V.5 The meaning of the horizontal error bars?
The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large. If all neutrinos were emitted simultaneously (a zero width time window) the recorded time of each neutrino would accurately represent its emission time, and there would be no need for having any horizontal error bars. If the time window were an amount $\Delta t$ then each neutrino might have its arriival time randomly shifted by a time up to that amount, and a horizonal error bar of that amount or greater would give a good fit to three straight lines. Thus, we can find the size of the emission time window by seeing for what horizontal error bar width H, the goodness of the fit probability significantly improves. Clearly from Fig. 2 it is about 1-2 sec.
V.6 Does a $\Delta t=1-2$ sec agree with simulations?
We can find the actual size of the time window using a simulation of neutrino emission. In 2025 Bozza et al. have calculated the neutrino luminosity from SN 1987A using a new model they recently developed. (Bozza2025, ) They claim the analysis is in many ways more refined and accurate than previous ones with respect to the temporal structure of the neutrino luminosity. The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig. 3. Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second. Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background. These results all buttress our case that the SN 1987A neutrinos were mostly emitted during a 1-2 sec time window, and that our use of 2 sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified. A 1 to 2 sec time window is apparently small enough not to badly degrade the fit of 30 neutrino data points to three straight lines.
V.7 Why don’t late-emitted $\nu$’s mess up the fit?
First, because based on Fig.3 there are not very many of these, perhaps no more than 15-20% of the total for $t>2$ sec, and second because of the late neutrinos associated with the shallower sloped line in Fig. 1, a shift of the points horizontally even by quite a few seconds on a shallow sloped line has little effect on the goodness of its fit to the line, which depends mainly on the vertical error bar. Thus, for example, a data point lying one sigma away vertically from that line could still be less than one sigma away if the point was shifted as much as five seconds horizontally.
Figure 3: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al. (Bozza2025, ). Note the different time scales used for $t<1$ sec and $t>1$ sec. The fraction of the integrated luminosity for $t<1$ is less than it seems because of this scale change. Figure is used with permission of publisher
V.8 Two really big vertical error bars?
Two data points occurring at the largest t in Fig. 1 appear anomalous particularly because of their very large error bars compared to all others. This is the result of error propagation and the fact that on a plot of $1/E^{2}$ versus t the size of the vertical error bar varies as the inverse cube of the energy E according to $\Delta(1/E^{2})=2\Delta E/E^{3}.$ In any case these two points degrade the fit very little because of their large error bars.
V.9 What if all three $m_{k}$ are all very small?
It is natural to wonder what Fig. 1 would look like if the three neutrino masses were very small and none were tachyonic, as in the standard neutrino description. Clearly, they still would lie on three lines, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics. However, we almost certainly would not be able to find these lines because for 3 small masses we would find three nearby steeply sloped lines, as in Fig.4. showing how the 3 lines might appear for neutrinos emitted at $t=0$ and $t=1$ sec Of course, with a horizontal error bar of 1 or 2 sec, identifying which of 30 points went with which nearly vertical line would be impossible because there are in fact many more line-triplets between $t=0$ and $t=1sec.$ In the “double bang” scenario for SN 1987A which could conceivably explain the LSD burst. However, once again we would not be able to identify the three lines about which the 30 points are clustering, as long as we are assuming very small neutrino masses. This means that the double bang scenario can be ruled out, given Fig.1.
Figure 4: If the 3 neutrino masses were very small as in the standard neutrino model, and none had $m_{j}^{2}<0,$ the three lines on $1/E^{2}$ versus t plot about which the 30 data points would cluster might look like Fig. 4, at least for those neutrinos emitted exactly at two times one second apart. But since neutrinos are in fact emitted at a continuum of times from $t=0$ to $t=1$ sec, there are many such line triplets will be the lines about which data points would cluster. Given 1 or 2 sec horizontal error bars, it clearly would be impossible to find any clustering of points about 3 specific lines, as is actually the case in Fig.1.
V.10 The preposterous size of the three $m^{2}_{j}$
The three masses deduced from the three straight line fit in Fig. 1 seem preposterously large, and they would appear to be inconsistent with data, whether it be comological data or data from the KATRIN experiment implying $m_{\nu}<1eV/c^{2}$. As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the effective mass has $m_{k}^{2}<0$ the effective mass can be arbitrarily close to zero or even negative. In order to look for the separate $m_{k}^{2}$ in a direct mass experiment like KATRIN one needs to combine three spectra one for each mass with appopriate weights (Ehrlich2019a, ). Likewise, the constraint imposed on the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ from cosmological data can be avoided if one or more of the tachyon masses is negative gravitationaly. (Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, so that they have a repulsive gravitational effect. The best upper limit on the total sum of neutrino masses from cosmology is $\sum m_{\nu}<0.12eV/c^{2}$ using data from the Planck satellite and the Atacama Cosmology Telescope. Without allowing negative $m_{k}$ all three would need to obey $m_{k}<0.12eV/c^{2},$ a limit that does not apply if negative $m_{k}$ are allowed. In light of these considerations, the numerical values from the three mass fit in Fig.1 are not necessarily ruled out by either lab or cosmological data.
V.11 The tail of the neutrino emission spectrum
It was noted earlier that background data not associated with the SN 1987A burst would not be expected to lie on three straight lines. Fortunately, Bozza et al. has obtained four Kamiokande II data points having times $15>t>25$ seconds on the day of SN 1989A. According to the “tail” of the emissions spectrum in Fig. 3, such data would have a negligible chance of being part of the SN 1987A burst, and they are likely to be background. Background events would be significantly lower in energy than supernova events. Because of that energy difference those four data points for $15>t>25$ seconds in fact have only a $1\%$ chance of being consistent with that best fit in Fig. 1 according to chi square.
VI what does it all mean?
The eleven considerations of the previous section suggest that the three neutrino masses found in the three line fit (Fig. 1) are real, even if one of the masses has $m^{2}<0.$ This is not the first time the author has claimed that the three neutrino masses can be found from SN 1987A data. When the evidence was first published in 2013, the three neutrino masses were interpreted in terms of a $3+3$ model, i.e., three active-sterile mass doublets, where each doublet was split by one of three oscillation $\Delta m^{2}.$ (Ehrlich2013, ; Ehrlich2019, ) We now discuss a much more plausible explanation involving CPT violation.
VI.1 Possible Violation of CPT Symmetry
CPT symmetry is at the heart of the standard model and is well-tested by looking for measured particle-antiparticle mass differences. CPT violation upper limits have been found for a number of pairs of particles and their antiparticles including proton-antiproton, electron-positron and $K_{0}-\bar{K}_{0}.$ Thus, for example $|m(K_{0})-m(\bar{K}_{0})|<0.50\times 10^{-18}GeV,$ which can also be expressed in dimensionless terms:
$X_{CPT}=\frac{|m(K_{0})-m(\bar{K}_{0})|}{m(K_{0})}<10^{-18}$ (6)
However, if viewed as a constraint on the mass-squared separation, the bound appears rather weak for kaons, namely only $0.25eV^{2}.$ It is generally believed that neutrinos might offer the best opportunity of any particle for observing CPT violation if it exists. The usual way of looking for violations or setting upper limits on CPT violation for neutrinos has not been to find a neutrino antineutrino mass difference, but rather to compare oscillation parameters for neutrinos and antineutrinos. There are currently two accelerator-based long baseline experiments T2K-II, NOvA-II and a reactor based medium-baseline JUNO experiment capable of testing CPT symmetry in this manner. The synergy of these three experiments should be able to set an upper bound in the $\Delta m^{2}$ between neutrinos and antineutrinos of $5.3\times 10^{-5}eV^{2}$, extending the current bound twenty fold. (Ngoc2023, ) Furthermore, an analysis of data from short-baseline experiments, combining Gallium radioactive source (neutrino) data and reactor (antineutrino) data, pointed to a possible CPT-violating asymmetry in the effective mixing angles with a statistical significance of about $3.5\sigma.$
The strongest hint of evidence for CPT violation has been provided by Hisakazu Minakata who notes that accelerator neutrino experiments in the $\nu_{\mu}\rightarrow\nu_{e}$ channel and the $\nu_{\mu}\rightarrow\nu_{\mu}$ channel show the problem of “appearance-disappearance tension” at a high confidence level. (Minakata2025, ) Minakata suggests non-unitarity as a simple and natural way of resolving the tension. He has constructed a model with a unique solution that is consistent with data, and implies a large neutrino-antineutrino asymmetry (hence CPT violation). Minakata’s model bridges the two highest CL signatures of CPT violation, i.e. data from the BEST and LSND-MiniBooNE experiments, and the model implies a large neutrino-antineutrino asymmetry.
As the preceding duscussion suggests, neutrinos are the ideal place to look for CPT violation because
  • Strong hints of CPT violation have been seen
  • The extent of the violation would be larger than for any other particles
  • The mechanism behind the tiny neutrino masses is not fully explained by the Standard Model.
  • Unlike charged particles, the weak neutrino interactions mean experiments are less susceptible to ”fake” CPT-violating signals that can arise from electromagnetic interactions.
  • Neutrinos can travel vast distances through matter without interacting significantly. Experiments use this to their advantage, studying neutrinos over very long baselines and at high energies.
A CPT violation might involve a mass difference between neutrino and antineutrino, which would bring about $\nu-\bar{\nu}$ oscillations. A relevant parameter specifying the magnitude of any CPT violation here is $(m^{2}_{\nu}-m^{2}_{\bar{\nu}})/m^{2}_{\nu}=\Delta m^{2}/m^{2}.$ Since there are three neutrino mass states one would expect that the CPT symmetry breaking parameter $X_{CPT}=\Delta m^{2}/m^{2}$ would be the same for the three masses, i.e.,
$X_{CPT}=\Delta m_{1}^{2}/m_{1}^{2}=\Delta m_{2}^{2}/m_{2}^{2}=\Delta m_{3}^{2}/m_{3}^{2}$ (7)
[…]
$X_{CPT}=\Delta m_{atm}^{2}/m_{1}^{2}=\Delta m_{sol}^{2}/m_{2}^{2}=\Delta m_{sbl}^{2}/m_{3}^{2}$ (8)
where we have used for $m_{1},m_{2}$ and $m_{3}$ in Eq. 7 the three masses found in our three mass fit to the SN 1987A data. Thus, we have the following dimensionless result for the amount of CPT violation from Eq. 8: $X_{CPT}=5\times 10^{-6},$ for the three mass states. Note that this postulated value is an order of magnitude smaller than that attainable in the three experiments T2K-II, NOvA-II and JUNO (which explains why no CPT violation has yet been directly seen).
The possible CPT violation represented by Eq. 8 may follow from a suggestion by Eichorn and Shiffer’s involving quantum gravity. (Eichorn2025, ) In the early universe during the transplankian regime when $\nu-\bar{\nu}$ oscillations did take place and the neutrino masses and their separations became frozen in. There is another implication of CPT violation for neutrinos involving tachyons. Various theoretical models of quantum gravity suggest it could lead to both CPT violation and modifications to the speed of neutrinos, resulting in subluminal and superluminal propagation, and so one might expect one or more member of the mass triplet defined by Eq. 5 be tachyonic. i.e., have $m^{2}_{j}<0.$ The recognition that neutrinos might provide evidence for CPT violation goes back as far as Bruno Pontecorvo (Pontecorvo1957, ) in 1957 who first suggested $\nu\bar{\nu}$ oscillations which can only occur if they have unequal masses.
VI.2 Why is CPT violation more likely than $3+3$?
The author previously postulated a $3+3$ model (3 active-sterile neutrino pairs) to explain the three outlandish masses found in the fit to the SN 1987A data (Ehrlich2019, ). There are many advantages of the CPT violation explanation over the earlier $3+3$ model. Thus, we have seen neutrinos are a natural place for CPT violation to occur, strong hints it is occurring, and there are difficulties with sterile neutrinos. The primary difficulty with sterile neutrinos is the contradiction with results from numerous other highly sensitive experiments which have found no evidence for them, as well as tension with cosmological data. In addition, with three sterile neutrinos and oscillations allowed between sterile and active neutrinos in the $3+3$ model one has a $6\times 6$ PMNS mixing matrix which appears to be in conflict with the standard well-established $3\times 3$ matrix parameters that are in good agreement with data.
Figure 5: An artists work depicting a 3D reconstruction based on SN 1987A observations. Credit ESO/L Calcada
VI.3 Summary and a conclusive test
This paper has made seven claims that some readers may consider very difficult to believe – see Table 1. Some of the seven claims will likely not be accepted as true without the strongest corroborating evidence.
1. No valid basis exists to ignore the LSD 5-hr early burst
2. No a priori upper limits exist on the three neutrino $m^{2}_{\nu,j}$
3. 3 $m_{j}^{2}$’s really can be found from 30 SN 1987A neutrinos
4. Negative search results cannot rule out tachyon’s existence
5. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
6. The three $m^{2}=2.7^{2}eV^{2},22.4^{2}eV^{2}$ and $2\times 10^{5}eV^{2}$
7. A likely explanation for these claims is CPT violation
Table 1: The seven claims made in this paper are based on the SN 1987A neutrino data. Regarding number 4, not only are negative search results incapable of proving tachyonic neutrinos non-existent, but they do not even necessarily make them less likely, but they may narrow the range of possible negative $m^{2}$ values.
We know SN 1987A was an asymmetric supernova (see Fig.5), with its expanding envelope and internal composition being lopsided rather than spherical. This spatial asymmetry was observed in the shape of the expanding envelope, and various other data. It is believed the asymmetry depicted likely arose from the progenitor star being a binary.
Figure 6: The asymmetric distribution of the debris of an explosion of the core of a 23-solar-mass progenitor star at 4.3 seconds after core bounce. The coloring codes mass density, blue being higher density. Credit: Argonne National Lab
Another particularly beautiful simulation of a supernova, this one lacking any symmetry whatsoever is shown in Fig. 6. However, it is not a mere spatial asymmetry that might corroborate the claims made in this paper, but rather a much deeper symmetry violation: the breaking of CPT symmetry and the evidence we have found supporting that violation by a dimensionless amount $X_{CPT}=5\times 10^{-5}.$ Usually, violations of the fundamental symmetries involve a decisive experiment. Thus, for parity (P) this was the C.S. Wu experiment, (Wu1957, ) for CP symmetry, this was the Fitch-Cronin experiment. (Fitch1964, ) Should a future experiment demonstrate a breaking of CPT symmetry for neutrinos by an amount $X_{CPT}=5\times 10^{-5}$ that would certainly be very strong corroboration of the seven claims. However, to find truly unambiguous evidence for all the claims, there is only one (and likely only one) decisive test, namely a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors.
The following concluding remarks represent one 88-year old’s supremely absurd optimism. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other six claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos, will definitely either confirm or reject the seven claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way to Earth from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say only 300 years). Were Betelgeuse to become a supernova, it would become incredibly bright, and even visible in daylight for many months. Go Betegeuse!
Acknowledgments
The author is grateful to Alan Chodos and Donald Gelman for helpful comments.
The author would be incredibly grateful to any readers
of this paper who find errors or poorly explained sections.
Conflicts of interest
The author declares no conflicts of interest.

### v2

@@
 Currently, the values of the three masses only have upper limits from cosmology and experiments like KATRIN.
-This paper shows that the SN 1987A neutrino data can remrkably yield values for the three neutrino masses, and not merely upper limits.
-Although this seemingly preposterous idea was suggested a dozen years ago by the author, here it is demonstrated in a much more convincing manner with many new elements, including a stronger statistical treatment, a theoretical linkage to possible CPT violation, and most importantly, a thorough explanation of why the method used to find the three masses from supernova SN 1987A neutrino data really works.
-The key to finding the three neutrino masses is realizing why three normally accepted assumotions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any masses $1eV/c^{2}$ would be inconsistent with upper limits from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A is too great for the method to work.
-Keywords: SN 1987A, Neutrino, supernova, CPT violation, tachyon, KATRIN
+This paper shows that the SN 1987A neutrino data can remarkably yield values for the three neutrino masses, and not merely upper limits.
+Although this seemingly preposterous idea was suggested a dozen years ago by the author, here it is demonstrated in a much more convincing manner with many new elements, including a stronger statistical treatment, and most importantly, a thorough explanation of why the method used to find the three masses from supernova SN 1987A neutrino data really works.
+The key to finding the three neutrino masses is realizing why three normally accepted assumptions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any neutrino masses $m_{k}>1eV/c^{2}$ cannot be reconciled with upper limits on the “effective mass” from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A data is too great for the method to work.
+A particularly interesting piece of evidence supporting the claim made in the paper’s title involves a recent negative KATRIN result finding an absence of sterile neutrinos.
+Keywords: SN 1987A, Neutrino, supernova, violation, tachyon, KATRIN, sterile neutrino
 I Introduction
@@
 Yet, thousands of papers have been published about SN 1987A, and its well-studied nature is not surprising given that it is the only supernova in our galaxy since neutrino detectors first existed.
-The early articles on the neutrino mass obtained from SN 1987A data mostly, with one exception,(Cowsik1988, ) set upper limits on its value.
-(Arnett1987, ; Burrows1988, ) The upper limits restriction is based on the almost universal belief that the differences in the three neutrino mass states are simply too small to use the data to find the individual masses themselves.
+The earliest articles on the neutrino mass obtained from SN 1987A data, with one exception,(Cowsik1988, ) set only upper limits on its value.
+(Arnett1987, ; Burrows1988, ; Fang1987, ).
+The upper limits restriction is based on the almost universal belief that the values of the three neutrino mass states are simply too small to use the SN 1987A data to find them.
 Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found.
-As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) an extensive discussion of CPT violation, (2) a much improved statistical analysis, (3) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A.
+As a result, the author rediscovered what Ramanath Cowsik had found in 1988 about the neutrino masses being possibly revealed in the SN 1987A data.
+(Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) a clue from the KATRIN experiment’s reported failure to observe sterile neutrinos.
+This absence surprisingly strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation of the preposterous-sounding claim.
+(KATRIN2025, )
 I.1 Are $m^{2}<0,$ and $v>c$ Tachyons “unphysical”?
@@
 Nevertheless, there are also theorists who are accepting of physical tachyons, and recently Paczos et al.
-(Paczos2024, ) have shown that a covariant framework exists which allows for the proper quantization of tachyon fields which eliminates all such “unphysical” issues.
-Charles Schwartz is another theorist who has challenged all the reasons why tachyons are considered unphysical.
-(Schwartz2022, ).
-I.2 The meaning of negative searches for tachyons
-The question of whether $m^{2}<0$ and $v>c$ observable particles really exist is not theoretical, but strictly empirical.
-All reports of tachyons, including the 2011 OPERA experiment initial $v>c$ result, (Brumfiel2012, ) have later been shown to be incorrect.
+(Paczos2024, ; Paczos2025, ) have shown that a covariant framework exists which allows for the proper quantization of tachyon fields which eliminates all such “unphysical” issues.
+Charles Schwartz is another theorist who has successfully challenged all the reasons why tachyons are considered unphysical, and shown how tachyons can account for many cosmological observations.
+(Schwartz2022, ; Schwartz2023, ).
+I.2 Do negative searches make tachyons less likely?
+The question of whether $m^{2}<0$ and $v>c$ observable particles exist is not theoretical, but strictly empirical.
+All reports of tachyons, including the 2011 OPERA experiment initial $v>c$ result, (Brumfiel2012, ) have later been shown to be incorrect apart from one-time sightings.
 (Ehrlich2022, ) Most physicists therefore believe that the tachyon hypothesis can be dismissed, or at least relegated to the highly implausible category.
@@
 A decline in belief in physical tachyons is perfectly understandable as a psychological reaction to negative results or even worse, false positive results like OPERA.
-However, negative and false positive results do not suggest tachyons do not exist, only that the $m^{2}$ for neutrinos is very close to zero, so we cannot tell if $m^{2}>0$ or $m^{2}<0$.
+However, negative and false positive results do not suggest tachyons do not exist, only that the $m^{2}$ for neutrinos is so close to zero, that we cannot tell if $m^{2}>0$ or $m^{2}<0$.
+You might think a negative hypothetical result $m^{2}=0.0\pm 0.1eV^{2}/c^{4}$ makes tachyonic neutrinos more unlikely than a previous result $m^{2}=0.0\pm 10.0eV^{2}/c^{4}$ because the allowed negative range in $m^{2}$ has been reduced a hundredfold.
+However, a Bayesian would realize that conclusion is false, because the allowed positive range has also been reduced by the same hundredfold amount.
 Neutrinos of course are the only known particles that might be tachyons since all others have been observed travelling at speeds $v<c,$ and tachyons if they exist must always be superluminal.
-Negative results for tachyonic neutrinos are not unlike negative results for searches for extraterestrials (SETI).
-In both cases, an appropriate response to a negative search would be “keep looking using more sensitive techniques.” Surprisinfly, however, if you ask an AI whether negative searches for tachyons make it more likely they don’t exist and then ask the same question about ET’s existence, you are illogically told yes for the former question and no for the latter.
-I.3 The KATRIN experiment
+I.3 KATRIN and circumventing upper limits on $m_{\nu}.$
 The KATRIN experiment seeks to find the “effective mass” of the electron antineutrino $m^{2}=\Sigma|U_{i,j}|^{2}m_{j}^{2}.$ Its best value as of 2025 is negative: $m^{2}=-0.14^{+0.13}_{-0.15}eV^{2}.$ (KATRIN2024, ) This result, of course, leaves the true sign of $m^{2}$ ambiguous given the size of the uncertainty.
-Is it just a coincidence that the KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass value the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}?$ Ehrlich2015 Maybe it is a just a weird coincidence, but fortunately, as KATRIN takes more data the statistical uncertainty on the $m^{2}$ value now $0.108eV^{2}$ will be reduced.
+Is it just a coincidence that the KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass value the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}?$ Ehrlich2015 Maybe it is a just a weird coincidence, but fortunately, as KATRIN takes more data the statistical uncertainty on the $m^{2}$ value will be reduced.
 Whatever the final KATRIN results, recall that the experiment measures the effective mass of the electron neutrino and not the three individual masses comprising it, which is the primary interest of this paper.
-II 3 Neutrino masses and SN 1987A
-It is widely believed that we can set an upper limit on the neutrino masses based on oscillation experiments.
-These experiments yield values for separations in the mass squared of pairs of mass states: $\Delta m^{2}_{j,k}=m_{j}^{2}-m_{k}^{2},$ the two well-established ones being the atmospheric and solar $m^{2}$ splittings $\Delta m^{2}_{atm}$ and $\Delta m^{2}_{sol}.$
-II.1 A true upper limit on Neutrino masses?
-The three masses $m^{2}_{k}$ values are normally assumed to be no greater than $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}\approx 0.0025eV^{2}.$ It is important to recognize this upper limit is based on a questionable assumption, namely the non-existence of any third observed oscillation frequency.
-In other words, since the two oscillation frequencies found for (1) $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}$ and (2) $\Delta m^{2}_{atm}$ are so close, they are empirically indistinguishable.
-If a third observed oscillation frequency did exist (perhaps the result of a fourth sterile neutrino) then the assumption of the upper limit $m_{k}^{2}<0.0025eV^{2}$ would be false.
-In that case, it is conceivable the neutrino masses are large enough to be measurable based on the neutrino travel time over a very large distance.
-Moreover, the best place to measure the three masses based on travel time might be a galactic supernova, provided the neutrino emissions from it are nearly simultaneous.
-No other source of a neutrino burst, even that produced by an atomic bomb, could come close to the precision made possible by the enormous travel distance to a supernova.
-II.2 Supernova SN 1987A neutrinos
+Later we shall see how upper limits on the three neutrino masses from KATRIN and cosmological data can be circumvented.
+Without any specific a priori upper limits, it is conceivable the neutrino masses are large enough to be measurable based on the neutrino travel time over a very large distance.
+The best baseline to measure the three masses based on their travel time is almost certainly a galactic supernova, provided we are confident the neutrino emissions from it are nearly simultaneous.
+An extragalactic one might seem even better were it not for the inverse square law and the very few neutrinos that could reach us.
+The phrase “neutrino burst” is used throughout this paper, referring to a sudden, intense emission of a vast number of neutrinos from a powerful cosmic event, most notably a core-collapse supernova lasting seconds.
+However we also use the phrase to refer to a seconds-long grouping of many more neutrinos than normally seen in a neutrino detector in such a time interval, which are believed to have originated from a supernova.
+I.4 Supernova SN 1987A neutrinos
 Three neutrino detectors operated at the time supernova SN1987A occurred at 7:35 UT on Feb 23, 1987.
@@
 The three detectors were not perfectly synchronized, with a time difference between them being perhaps a few seconds.
-This lack of synch is partially removed in most analyses by assuming the first event in each detector occurred at time $t=0.$ The arriving electron antineutrinos were detected based on inverse beta decay
+This lack of synchronization is usually accounted for by assuming the first event in each detector occurred at time $t=0.$ The arriving electron antineutrinos were detected based on inverse beta decay
 $\bar{\nu_{e}}+p\rightarrow n+e^{+}$ (1)
@@
 Those three neutrino bursts occurred about three hours before the star emitting them suddenly brightened.
-This late light arrival, of course does not imply that neutrinos are superlumnial.
+This late light arrival, of course does not imply that neutrinos are superluminal.
 Rather, the three-hour gap between the neutrino burst and the later arrival of visible light, aligns with the small interaction between neutrinos with matter.
-Thus, photons but not neutrinos are delayed by interactions with the supenova ejecta.
-II.3 The 5 hour early LSD neutrino burst
+Thus, photons but not neutrinos are delayed by interactions with the supernova ejecta.
+I.5 The 5 hour early LSD neutrino burst
 In addition to the three detectors, there was a fourth small detector known as LSD ln 1987 which was located under Mont Blanc.
-LSD reported a burst of five neutrinos which occurred 484 min before the other three detectors’ bursts.
+This detector was in fact better shielded from background radiation than the other three, and had a lower energy threshold, so in these two respects was more sensitive.
+The LSD detector reported a burst of five neutrinos which occurred 484 min before the other three detectors’ received bursts.
 This early burst is therefore considered by most physicists to be a background fluctuation unrelated to SN 1987A because of its five hour early arrival.
-If the LSD burst were genuine it could constitute evidence that those neutrinos were tachyons, although other explanations also exist, including having SN1987A “bang” twice (Galeotti2016, ) and other ones ones as well.
+If the LSD burst were genuine it could constitute evidence that those neutrinos were tachyons, although other explanations also exist, including having SN1987A “bang” twice (Galeotti2016, ) and other ones as well.
 (Agafonova2018, ) Nevertheless, even raising the possibility of tachyons probably was enough to dismiss the existence of the LSD neutrino burst as unrelated to SN 1987A in the minds of most physicists.
-The five neutrinos observed during a 7 second time interval in LSD would if spurious represent a very significant departure from background (roughly one neutrino every 80 seconds), and that background had been steady for months.
+On the other hand, there were four other reasons to be suspicious about the LSD burst besides the 5 hr early arrival, starting with it possibly being a statistical fluctuation of background.
+The five neutrinos observed during a 7 second time interval in LSD would, if not connected to SN1987A, represent a very significant departure from background (roughly one neutrino every 80 seconds), and that background had been steady for months.
 In fact, the probability of a chance fluctuation giving rise to the LSD signal is extremely low, less than $1.4\times 10^{–6}.$ (Galeotti2016, ) Many observers also reject the LSD burst because no such early burst was seen in the three other detectors.
@@
 However, that can also be explained because the small LSD detector did observe two neutrinos that time (not enough to qualify as a burst).
-Two final major problem with the LSD burst being associated with SN 1987A are (1) why the very low energies of all the 5 neutrinos comprising it ?, and (2 )why the virtual monochromaticity of the burst?
-In what follows we accept that the LSD neutrino burst was associated with SN 1987A, because all five above objections to it are answerable, and we address the monochromaticity problem later.
-III Finding $m^{2}$ of SN 1987A neutrinos
-The possibility of using the SN 1987A data to find the three neutrino masses may seem implausible, but we shall see that it can be done if the neutrinos were emitted from the star close to simultaneously.
-Let us first see why for exactly simultaneous neutrino emissions the neutrino arrival time relative to a photon directly determines their travel time (relative to light), and how using the measured neutrino energy $E,$ and arrival time t we can compute the mass of $\emph{i}ndividual$ neutrinos.
+Two final major problem with the LSD burst being associated with SN 1987A are (1) the very low energies of all the 5 neutrinos comprising it, and (2 ) their nearly equal energies or the virtual monochromaticity of the burst.
+In what follows we accept that the LSD neutrino burst was associated with SN 1987A, because all five objections to it are answerable, but we address the monochromaticity problem later.
+II Finding $m^{2}$ of SN 1987A neutrinos
+The possibility of using the SN 1987A data to find the three neutrino masses may seem implausible, but we shall see that it can be done if the neutrinos were emitted from the star close enough to simultaneously.
+Let us first see why for exactly simultaneous neutrino emissions the neutrino arrival time relative to a photon directly determines their travel time (relative to light), and how using the measured neutrino energy $E,$ and arrival time t we can compute the mass of the neutrinos.
 In what follows $T=168,000$ years is the time photons required to reach us from SN 1987A.
@@
 For the neutrino speed we have v/c=1t/T.v/c=1-t/T.
-and also $v/c=\sqrt{1-m^{2}c^{4}/E^{2}}\approx 1-m^{2}c^{4}/2E^{2}.$ Setting the two versions of $v/c$ equal
+and also as an excellent approximation $v/c=\sqrt{1-m^{2}c^{4}/E^{2}}\approx 1-m^{2}c^{4}/2E^{2}$.
+Setting the two versions of $v/c$ equal
 $v/c=1-t/T=1-m^{2}c^{4}/2E^{2}$ (2)
@@
 Eq.
-3 implies that if a number of neutrinos with the same mass $m^{2}c^{4}$ were present and they had different arrival times t and energies E, then on a plot of $1/E^{2}$ versus $t,$ the neutrinos will be found to lie on or near a straight line through the origin of slope M given by
+3 implies that if a number of neutrinos with the same mass $m^{2}c^{4}$ were present and they had different arrival times t and energies E, then on a plot of $1/E^{2}$ versus $t,$ the data point for each neutrino will be found to lie on or near a straight line through the origin of slope M given by
 $M=\frac{2}{Tm^{2}c^{4}}$ (4)
@@
 $m^{2}c^{4}=\frac{2}{TM}$ (5)
-Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line having negative slope.
-To the extent that neutrinos all started out from SN 1987A at nearly the same instant, then since there are known to be three (and only three) neutrino masses, every neutrino should lie on or close to one of three straight lines passing through the origin.
+Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line through the origin having negative slope.
+To the extent that neutrinos all started out from SN 1987A at nearly the same instant, then since there are known to be three (and only three) non-sterile neutrino masses, every neutrino should lie on or close to one of three straight lines passing through the origin.
 Any neutrinos that do not do so must either be a background event unrelated to SN 1987A, or else the result of non-simultaneous emissions.
-It is however possible that the data points will all be found to be associated with only one or two straight lines rather than three lines if two or three of the neutrino masses are too close to distinguish from one another.
-Before we examine what the data show, there is a philosophical issue to mention.
-Given that in the standard view neutrino masses are less than $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}$ – a value so small that any spread in neutrino travel times could not possibly be large enough to observe the neutrino mass, we have two choices on how to proceed.
-The first is not to even bother to see what the data suggest (the ”Cremonini option”) given that we know any positive result must be an anomaly, given the small upper limt on the neutrino mass .
-However, we have made the second choice, namely not to make any assumptions on the size of the neutrino masses and be guided by what the data themselves show.
-IV Fitting 30 data points to 3 lines
-Strikingly, when the SN 1987A data is plotted (see Fig.1) every one of the 30 neutrinos lies on one of three straight lines through the origin, and very few are next to two lines.
-The least square fit of the 30 data points to three lines through the origin has two free parameters.
-The first parameter is $S$ which adjusts the trial slopes of the three straight lines in Fig.
-1 simultaneously and the second free parameter H is the size of the constant horizontal error bar on each data point in Fig.1.
+It is however possible that the data points will all be found to be associated with only one or two straight lines rather than three lines if two or three of the lines (or neutrino masses) are too close to distinguish from one another.
+III Fitting 30 data points to 3 lines
+Thirty neutrinos is not a very large number, but the distinctiveness of a pattern counts for far more than the sheer number of events.
+Thus, the $\Omega^{-}$ particle which led to quarks was confirmed in 1964 at Brookhaven National Laboratory by observing just one single“Gold Standard” event.
+(Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the 30 neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only three points near the origin are unavoidably next to two lines.
+The least square fit of the 30 data points to three lines through the origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig.
+1 yielding 27 degrees of freedom.
+The parameter H, the size of the constant horizontal error bar on each data point in Fig.1 is also adjusted.
 Since each data point has both x and y error bars, we can then convert the x error into an ”effective” y error by multiplying the x uncertainty by the slope of the line, and combine it with the y uncertainty in quadrature.
 If a data point lies close to two straight lines, we simply choose the line that gives the smaller contribution to chi square.
-However, only a few of the data points near the origin are within $1\sigma$ of more than one straight line.
+Only a few of the data points lie beyond $1\sigma$ of a straight line.
 The good fit shown in Fig 1 has a chi square of 25.4.
-Given 28 degrees of freedom we find $p=61\%.$ The excellent best fit was for a horizontal error bar of two seconds on all data points, and the fit would still be acceptable but significantly worse for a one second horizontal error bar ($p=18\%$.) For a fuller desccription of how the $\chi^{2}$ fit probability depends on the size of the horizontal error bar see Fig.
+Given 27 degrees of freedom we find $p=55\%.$ This excellent best fit was for a horizontal error bar of two seconds on all data points, and the fit would still be acceptable but significantly worse for a one second horizontal error bar ($p=15\%$.) For a fuller description of how the $\chi^{2}$ fit probability depends on the size of the horizontal error bar see Fig.
 2 We have seen that exactly simultaneous neutrino emissions requires the data must lie on one of three lines through the origin in a $1/E^{2}$ versus t plot.
-Therefore the goodness of fits using 1 and 2 sec horizontal error bars shows they will continue to lie on or near these lines if the time window of neutrino emissions is no more than 1-2 seconds.
-Figure 1: Best fit of 30 data points $(t,1/E^{2}),$ one for each neutrino to three straight lines that pass through the origin.
+Therefore the goodness of fits using 1 and 2 sec horizontal error bars shows they will continue to lie on or near these lines if the time window of neutrino emissions is also no more than 1-2 seconds.
+Figure 1: Best fit of 30 data points $(t,1/E^{2}),$ one for each neutrino to one of three straight lines that pass through the origin.
 Error bars in both horizontal and vertical are included in the fit.
-The best fit has $\chi^{2}=25.0$ for 28 dof, which yields $p=61\%$.
+The fit has $\chi^{2}=25.4$ for 27 dof, which yields $p=55\%$.
 The three best fit masses $m_{j}$ are obtained from the three best fit slopes according to Eq.
@@
 Note the different scales used for $t>0$ and $t<0$ The five LSD data points at t=17ksect=-17ksec have been slightly separated from each other for clarity.
-The three masses corresponding to the slopes of the three straight lines are $m_{1}=22.4eV/c^{2}$ and $m_{2}=2.7eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ for the $m^{2}<0$ line.
-The big difference from the author’s previously published values of this fit is in the value of $m_{2}$ because of the past omission of horizontal error bars, which are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line.
-We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.7eV/c^{2}$ and the $2.7eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration, as well as its non-observation by KATRIN.
-(Serebrov2021, ; KATRIN2025, ).
-Of course, our $m_{1},m_{2},m_{3},$ however, are not sterile neutrino masses, but the three active neutrino masses.
+The meaning of the shaded central area is discussed in the text
+The three masses corresponding to the slopes of the three straight lines in Fig.
+1 are $m_{1}=22.4eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ for the $m^{2}<0$ line.
+The author first discovered this three line fit in 2013, when he interpreted the result in terms of a 3 + 3 model (three pairs of active stable neutrinos – see Fig.3.) The big difference from the author’s previously published values of this fit (Ehrlich2013, ; Ehrlich2019, ) is in the value of $m_{2}$ because of the past omission of horizontal error bars.
+Those error bars are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line.
+We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.70eV/c^{2}$ and the $2.70eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration (both equal to within $1\%$), as well as its non-observation by KATRIN, (Serebrov2024, ; KATRIN2025, ) a matter discussed at length in section IV.9.
+III.1 Fixing a limitation of sterile neutrino searches
+Of course, the existence of sterile neutrinos is highly controversial, there are probably more negative search results than positive ones.
+Thus, neither MicroBooNE, nor IceCube, nor KATRIN nor DayaBay report any evidence for them.
+(Micro2025, ; Aartsen2020, ; KATRIN2025, ; An2024, ).
+Nevertheless, an omission in these and other experiments is that they have not sought evidence for the presence of multiple sterile neutrinos simultaneously.
+In effect, if you are only searching for one sterile neutrino when two or three are present they can mask each others presence.
+One way to avoid this problem is to search for two or three in the same fit to the data, and another way is to find a way to spatially separate any sterile neutrinos based on their masses, or find a place where nature causes this separation for us.
+III.2 More evidence for the two $m^{2}>0$ masses
+It is believed that dark matter holds together both our galaxy as well as clusters of galaxies.
+One would expect that if both heavy and light $m^{2}>0$ sterile neutrinos were present throughout space the heavy ones would preferentially congregate in an individual galaxy and the lighter ones would preferentially congregate an a galaxy cluster.
+M.
+H.
+Chan and the author have carried out fits to the “observed” dark matter distributions in both the Milky Way galaxy and in four clusters of galaxies.
+(Chan2014, ) The goal was to see what mass sterile neutrino would match the observed profile in each case.
+This was done by deriving an equation of state for the sterile neutrinos dependent on their mass, and then finding the dark matter radial profile in both the Milky Way and four groups of galaxies.
+Finally, the mass was varied to find the value that best fit the observed profile in each case.
+The results of these calculations showed that the two $m^{2}>0$ values from the 3-mass SN 1987A fit would do nicely with the heavier mass holding clusters together as would be expected.
+It seems unlikely one could find another case where nature is going to be so kind as to separate spatially different mass dark matter particles.
+However, as we shall see in section IV.9, KATRIN has a way to search for all three neutrino masses including the tachyonic one simultaneously in a single fit.
 Figure 2: The probability from the least squares fit to three straight lines through the origin versus the size of the horizontal error bars, assumed to be the same for all 30 data points.
-Obviously for $H>>5$ sec error bars the fit has $p\approx 100\%$
-V Discussion
-In this section we discuss eleven issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0,$ really does imply three specifc neutrino masses, and it is not just a mirage or background fluctuation.
-V.1 Why not not exclude the LSD data?
-Initially, the author rejected the LSD data.
-This rejection was not because it proviided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous.
-However since this and all the other reasons many people have for rejecting the LSD data can be answered, there really is no valid reason not to keep it.
-V.2 Why are the LSD data bunched in time?
-If the LSD burst of 5 neutrinos is genuine, it looks strange in Fig.
+Note that there is no “best” fit to three straight lines unless one first fixes the size of the horizontal error bars because for $H>>5$ sec error bars the fit probability $p\approx 100\%$
+Figure 3: A possible $3+3$ neutrino mass squared diagram postulated by the author in 2013, showing updated $m_{1}$ and $m_{2}$ values.
+There are 3 active-sterile doublets (solid and dotted lines) split by three $\Delta m^{2}$ The three $\Delta m^{2}$ are for solar, atmospheric and short baseline (sbl) oscillations.
+Diagram is not to scale.
+The $m_{2}$ value differs significantly from the previously found value, i.e.
+$(m_{2}=4.0eV/c^{2})$ rather than $2.70eV/c^{2},$ because of the use of horizontal error bars now used in the three mass fit
+IV Discussion
+In this section we discuss 10 issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0,$ really does imply we can deduce the three neutrino masses from the SN 1987A data, and it is not just a mirage or background fluctuation.
+IV.1 Why not exclude the LSD data?
+Initially, the author mistakenly rejected the LSD data.
+This rejection was not because it provided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous.
+However since this issue and all the other reasons many people mistakenly have for rejecting the LSD data can be answered, there is no valid reason not to keep LSD data, especially because with three neutrinos the mandated number of straight lines about which clustering should occur is three.
+IV.2 Why are the LSD data so bunched in time?
+If the LSD burst of 5 neutrinos is genuine, it may look strange in Fig.
 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec.
-The three straight line result would be more believable if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line.
+The three straight line result would be more believable perhaps if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line.
 But recall that the $t<0$ data has a kilosecond time scale.
-If the five LSD data points were spread out along the negative sloped line and not bunched in time they never could have been distinguished from background because researchers were looking for a bunch of events in a seconds-long time interval.
-V.3 Why do the LSD neutrinos have the same E?
+If the five LSD data points were spread out along the negative sloped line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst.
+IV.3 Why do the LSD neutrinos have the same E?
 The apparent near-constant energy of the LSD burst remains mysterious even to the physicists who accept its existence, but it is natural if the 5 neutrinos (and all others) must lie on one of three straight lines.
 Moreover, there is a theoretical basis for the monochromaticity of the LSD burst, as shown in several of the author’s publications that have provided independent evidence for it.
-That evidence was in the form of a proposed model for an 8 MeV neutrino line from a supernova, and further evidence related to the observed spectrum of MeV $\gamma$-rays from the galactic center.
+That evidence was in the form of a proposed model for an 8 MeV neutrino line from a supernova, and support for the model from the observed spectrum of MeV $\gamma$-rays from the galactic center.
 (Ehrlich2021, )
-V.4 Alternate explanations for the LSD neutrinos?
-All the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems. Thus,
+IV.4 Alternate explanations for the LSD neutrinos?
+Many phenomena in nature on the fringes of the known have alternative explanations. All the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems. Thus,
   • If it was a statistical fluctuation it was a really big one $p=1.4\times 10^{-6}$
-  • If it was due to a double bang the pattern in Fig. 1 should look very different as discussed in section 6.
+  • If it was due to a double bang the pattern in Fig. 1 should look very different as discussed in section IV.7.
   • If it was a genuine, low-energy pulse (e.g., from neutron capture on iron slabs around the detector), it would require the supernova to release an unusually high amount of energy in this form.
   • If it was due to 8 MeV dark matter as suggested in ref.
-(Ehrlich2021, ) this would still creates the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD neutrinos falling on the $m^{2}<0$ line.
-V.5 The meaning of the horizontal error bars?
-The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large.
+(Ehrlich2021, ) this would still create the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD neutrinos falling on the $m^{2}<0$ line.
+IV.5 Best size of the horizontal error bars?
+The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large for the method to work.
 If all neutrinos were emitted simultaneously (a zero width time window) the recorded time of each neutrino would accurately represent its emission time, and there would be no need for having any horizontal error bars.
-If the time window were an amount $\Delta t$ then each neutrino might have its arriival time randomly shifted by a time up to that amount, and a horizonal error bar of that amount or greater would give a good fit to three straight lines.
+If the time window were an amount $\Delta t$ then each neutrino might have its arrival time t randomly shifted to the right by a time up to that amount, and a horizonal error bar $H\approx\Delta t$ would give a good fit to three straight lines.
 Thus, we can find the size of the emission time window by seeing for what horizontal error bar width H, the goodness of the fit probability significantly improves.
 Clearly from Fig.
-2 it is about 1-2 sec.
-V.6 Does a $\Delta t=1-2$ sec agree with simulations?
-We can find the actual size of the time window using a simulation of neutrino emission.
+2 it is about 1-2 seconds.
+IV.6 The time window for neutrino emissions?
+We can estimate the actual size of the time window for neutrino emissions using a simulation.
 In 2025 Bozza et al.
@@
 The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig.
-3.
+4.
 Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second.
@@
 A 1 to 2 sec time window is apparently small enough not to badly degrade the fit of 30 neutrino data points to three straight lines.
-V.7 Why don’t late-emitted $\nu$’s mess up the fit?
-First, because based on Fig.3 there are not very many of these, perhaps no more than 15-20% of the total for $t>2$ sec, and second because of the late neutrinos associated with the shallower sloped line in Fig.
-1, a shift of the points horizontally even by quite a few seconds on a shallow sloped line has little effect on the goodness of its fit to the line, which depends mainly on the vertical error bar.
-Thus, for example, a data point lying one sigma away vertically from that line could still be less than one sigma away if the point was shifted as much as five seconds horizontally.
 Figure 4: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al.
 (Bozza2025, ).
-Note the different time scales used for $t<1$ sec and $t>1$ sec.
-The fraction of the integrated luminosity for $t<1$ is less than it seems because of this scale change.
-Figure is used with permission of publisher
-V.8 Two really big vertical error bars?
-Two data points occurring at the largest t in Fig.
-1 appear anomalous particularly because of their very large error bars compared to all others.
-This is the result of error propagation and the fact that on a plot of $1/E^{2}$ versus t the size of the vertical error bar varies as the inverse cube of the energy E according to $\Delta(1/E^{2})=2\Delta E/E^{3}.$ In any case these two points degrade the fit very little because of their large error bars.
-V.9 What if all three $m_{k}$ are all very small?
+Note the time scale change at $t=1$ sec.
+The fraction of the integrated luminosity for $t<1$ is less than it seems because of the scale change.
+Figure is used under creative commons license https://creativecommons.org/licenses/by/4.0
+IV.7 Suppose neutrinos all have $m<<1eV/c^{2}$?
 It is natural to wonder what Fig.
@@
 However, we almost certainly would not be able to find these lines because for 3 small masses we would find three nearby steeply sloped lines, as in Fig.5.
-showing how the 3 lines might appear for neutrinos emitted at $t=0$ and $t=1$ sec Of course, with a horizontal error bar of 1 or 2 sec, identifying which of 30 points went with which nearly vertical line would be impossible because there are in fact many more line-triplets between $t=0$ and $t=1sec.$ In the “double bang” scenario for SN 1987A which could conceivably explain the LSD burst.
-However, once again we would not be able to identify the three lines about which the 30 points are clustering, as long as we are assuming very small neutrino masses.
+showing how the 3 lines might appear for neutrinos emitted at $t=0$ and $t=1$ sec Of course, with a horizontal error bar of 1 or 2 sec, identifying which of 30 points went with which nearly vertical line would be impossible because there are in fact many more line-triplets between $t=0$ and $t=1sec.$
+If the neutrino mass were in fact under the KATRIN current upper limit $(0.45eV/c^{)}$ the data points would all lie on lines that were virtually vertical and the data points would all lie in the shaded central area in Fig.1 having a 2 second time window.
+This is obviously far from reality, which suggests that their mass is not below the KATRIN limit.
+In the “double bang” scenario for SN 1987A which could conceivably explain the LSD burst, the preceding remains true, so once again we would not be able to identify the three lines about which the 30 points are clustering, unlike the actual data in Fig.
+1.
 This means that the double bang scenario can be ruled out, given Fig.1.
-Figure 4: If the 3 neutrino masses were very small as in the standard neutrino model, and none had $m_{j}^{2}<0,$ the three lines on $1/E^{2}$ versus t plot about which the 30 data points would cluster might look like Fig.
-4, at least for those neutrinos emitted exactly at two times one second apart.
-But since neutrinos are in fact emitted at a continuum of times from $t=0$ to $t=1$ sec, there are many such line triplets will be the lines about which data points would cluster.
-Given 1 or 2 sec horizontal error bars, it clearly would be impossible to find any clustering of points about 3 specific lines, as is actually the case in Fig.1.
-V.10 The preposterous size of the three $m^{2}_{j}$
+Figure 5: The case of neutrinos having very small masses (which yield triplets of steeply sloped straight lines nearly overlapping) with a data point having a 2 sec wide horizontal error bar.
+Neutrinos are assumed emitted at $t=0$ and $t=1$ seconds.
+Given the size of the error bar, it would be impossible to find which of the six near vertical lines the data point is associated with (likewise for all the other line triplets for $0>t>1$ seconds – completely unlike what is actually the case in Fig.1.
+IV.8 Consistency with direct mass experiments?
 The three masses deduced from the three straight line fit in Fig.
-1 seem preposterously large, and they would appear to be inconsistent with data, whether it be comological data or data from the KATRIN experiment implying $m_{\nu}<1eV/c^{2}$.
-As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the effective mass has $m_{k}^{2}<0$ the effective mass can be arbitrarily close to zero or even negative.
-In order to look for the separate $m_{k}^{2}$ in a direct mass experiment like KATRIN one needs to combine three spectra one for each mass with appopriate weights (Ehrlich2019a, ).
-Likewise, the constraint imposed on the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ from cosmological data can be avoided if one or more of the tachyon masses is negative gravitationaly.
-(Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, so that they have a repulsive gravitational effect.
-The best upper limit on the total sum of neutrino masses from cosmology is $\sum m_{\nu}<0.12eV/c^{2}$ using data from the Planck satellite and the Atacama Cosmology Telescope.
-Without allowing negative $m_{k}$ all three would need to obey $m_{k}<0.12eV/c^{2},$ a limit that does not apply if negative $m_{k}$ are allowed.
-In light of these considerations, the numerical values from the three mass fit in Fig.1 are not necessarily ruled out by either lab or cosmological data.
-V.11 The tail of the neutrino emission spectrum
-It was noted earlier that background data not associated with the SN 1987A burst would not be expected to lie on three straight lines.
-Fortunately, Bozza et al.
-has obtained four Kamiokande II data points having times $15>t>25$ seconds on the day of SN 1989A.
-According to the “tail” of the emissions spectrum in Fig.
-3, such data would have a negligible chance of being part of the SN 1987A burst, and they are likely to be background.
-Background events would be significantly lower in energy than supernova events.
-Because of that energy difference those four data points for $15>t>25$ seconds in fact have only a $1\%$ chance of being consistent with that best fit in Fig.
-1 according to chi square.
-VI what does it all mean?
-The eleven considerations of the previous section suggest that the three neutrino masses found in the three line fit (Fig.
-1) are real, even if one of the masses has $m^{2}<0.$ This is not the first time the author has claimed that the three neutrino masses can be found from SN 1987A data.
-When the evidence was first published in 2013, the three neutrino masses were interpreted in terms of a $3+3$ model, i.e., three active-sterile mass doublets, where each doublet was split by one of three oscillation $\Delta m^{2}.$ (Ehrlich2013, ; Ehrlich2019, ) We now discuss a much more plausible explanation involving CPT violation.
-VI.1 Possible Violation of CPT Symmetry
-CPT symmetry is at the heart of the standard model and is well-tested by looking for measured particle-antiparticle mass differences.
-CPT violation upper limits have been found for a number of pairs of particles and their antiparticles including proton-antiproton, electron-positron and $K_{0}-\bar{K}_{0}.$ Thus, for example $|m(K_{0})-m(\bar{K}_{0})|<0.50\times 10^{-18}GeV,$ which can also be expressed in dimensionless terms:
-$X_{CPT}=\frac{|m(K_{0})-m(\bar{K}_{0})|}{m(K_{0})}<10^{-18}$ (6)
-However, if viewed as a constraint on the mass-squared separation, the bound appears rather weak for kaons, namely only $0.25eV^{2}.$ It is generally believed that neutrinos might offer the best opportunity of any particle for observing CPT violation if it exists.
-The usual way of looking for violations or setting upper limits on CPT violation for neutrinos has not been to find a neutrino antineutrino mass difference, but rather to compare oscillation parameters for neutrinos and antineutrinos.
-There are currently two accelerator-based long baseline experiments T2K-II, NOvA-II and a reactor based medium-baseline JUNO experiment capable of testing CPT symmetry in this manner.
-The synergy of these three experiments should be able to set an upper bound in the $\Delta m^{2}$ between neutrinos and antineutrinos of $5.3\times 10^{-5}eV^{2}$, extending the current bound twenty fold.
-(Ngoc2023, ) Furthermore, an analysis of data from short-baseline experiments, combining Gallium radioactive source (neutrino) data and reactor (antineutrino) data, pointed to a possible CPT-violating asymmetry in the effective mixing angles with a statistical significance of about $3.5\sigma.$
-The strongest hint of evidence for CPT violation has been provided by Hisakazu Minakata who notes that accelerator neutrino experiments in the $\nu_{\mu}\rightarrow\nu_{e}$ channel and the $\nu_{\mu}\rightarrow\nu_{\mu}$ channel show the problem of “appearance-disappearance tension” at a high confidence level.
-(Minakata2025, ) Minakata suggests non-unitarity as a simple and natural way of resolving the tension.
-He has constructed a model with a unique solution that is consistent with data, and implies a large neutrino-antineutrino asymmetry (hence CPT violation).
-Minakata’s model bridges the two highest CL signatures of CPT violation, i.e.
-data from the BEST and LSND-MiniBooNE experiments, and the model implies a large neutrino-antineutrino asymmetry.
-As the preceding duscussion suggests, neutrinos are the ideal place to look for CPT violation because
-  • Strong hints of CPT violation have been seen
-  • The extent of the violation would be larger than for any other particles
-  • The mechanism behind the tiny neutrino masses is not fully explained by the Standard Model.
-  • Unlike charged particles, the weak neutrino interactions mean experiments are less susceptible to ”fake” CPT-violating signals that can arise from electromagnetic interactions.
-  • Neutrinos can travel vast distances through matter without interacting significantly. Experiments use this to their advantage, studying neutrinos over very long baselines and at high energies.
-A CPT violation might involve a mass difference between neutrino and antineutrino, which would bring about $\nu-\bar{\nu}$ oscillations.
-A relevant parameter specifying the magnitude of any CPT violation here is $(m^{2}_{\nu}-m^{2}_{\bar{\nu}})/m^{2}_{\nu}=\Delta m^{2}/m^{2}.$ Since there are three neutrino mass states one would expect that the CPT symmetry breaking parameter $X_{CPT}=\Delta m^{2}/m^{2}$ would be the same for the three masses, i.e.,
-$X_{CPT}=\Delta m_{1}^{2}/m_{1}^{2}=\Delta m_{2}^{2}/m_{2}^{2}=\Delta m_{3}^{2}/m_{3}^{2}$ (7)
-[…]
-$X_{CPT}=\Delta m_{atm}^{2}/m_{1}^{2}=\Delta m_{sol}^{2}/m_{2}^{2}=\Delta m_{sbl}^{2}/m_{3}^{2}$ (8)
-where we have used for $m_{1},m_{2}$ and $m_{3}$ in Eq.
-7 the three masses found in our three mass fit to the SN 1987A data.
-Thus, we have the following dimensionless result for the amount of CPT violation from Eq.
-8: $X_{CPT}=5\times 10^{-6},$ for the three mass states.
-Note that this postulated value is an order of magnitude smaller than that attainable in the three experiments T2K-II, NOvA-II and JUNO (which explains why no CPT violation has yet been directly seen).
-The possible CPT violation represented by Eq.
-8 may follow from a suggestion by Eichorn and Shiffer’s involving quantum gravity.
-(Eichorn2025, ) In the early universe during the transplankian regime when $\nu-\bar{\nu}$ oscillations did take place and the neutrino masses and their separations became frozen in.
-There is another implication of CPT violation for neutrinos involving tachyons.
-Various theoretical models of quantum gravity suggest it could lead to both CPT violation and modifications to the speed of neutrinos, resulting in subluminal and superluminal propagation, and so one might expect one or more member of the mass triplet defined by Eq.
-5 be tachyonic.
-i.e., have $m^{2}_{j}<0.$ The recognition that neutrinos might provide evidence for CPT violation goes back as far as Bruno Pontecorvo (Pontecorvo1957, ) in 1957 who first suggested $\nu\bar{\nu}$ oscillations which can only occur if they have unequal masses.
-VI.2 Why is CPT violation more likely than $3+3$?
-The author previously postulated a $3+3$ model (3 active-sterile neutrino pairs) to explain the three outlandish masses found in the fit to the SN 1987A data (Ehrlich2019, ).
-There are many advantages of the CPT violation explanation over the earlier $3+3$ model.
-Thus, we have seen neutrinos are a natural place for CPT violation to occur, strong hints it is occurring, and there are difficulties with sterile neutrinos.
-The primary difficulty with sterile neutrinos is the contradiction with results from numerous other highly sensitive experiments which have found no evidence for them, as well as tension with cosmological data.
-In addition, with three sterile neutrinos and oscillations allowed between sterile and active neutrinos in the $3+3$ model one has a $6\times 6$ PMNS mixing matrix which appears to be in conflict with the standard well-established $3\times 3$ matrix parameters that are in good agreement with data.
-Figure 5: An artists work depicting a 3D reconstruction based on SN 1987A observations. Credit ESO/L Calcada
-VI.3 Summary and a conclusive test
-This paper has made seven claims that some readers may consider very difficult to believe – see Table 1.
-Some of the seven claims will likely not be accepted as true without the strongest corroborating evidence.
+1 seem preposterously large, and they would appear to be inconsistent with data, whether it be cosmological data or data from the KATRIN experiment implying $m_{\nu}<0.45eV/c^{2}$.
+As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the electron neutrino effective mass has $m_{k}^{2}<0$ the effective mass, as normally defined, $m_{e}^{2}=\sum|U_{e,k}|^{2}m_{k}^{2}$ can be arbitrarily close to zero or even negative, even when the $m_{k}^{2}$ are all very large, as in the SN 1987 fit.
+KATRIN has further crucial relevance to that fit quite apart from the limit it imposes by its measured $\nu_{e}$ effective mass.
+IV.9 A missing $m_{4}=2.70eV/c^{2}$ sterile neutrino?
+KATRIN recently reported finding no evidence for sterile neutrinos in its data based on the observed shape of the tritium beta decay spectrum.
+(KATRIN2025, ).
+It conducted this search by fitting its data to a sum of two spectra: that for a zero mass neutrino plus the spectrum for a sterile neutrino of mass $m_{4}$, which was assumed to take on various masses and various weights.
+KATRIN then used a logarithmic grid of trial $m^{2}$ and $sin^{2}2\theta$ values that spans from $m^{2}=0.1eV^{2}$ to $1600eV^{2}$ and weight $sin^{2}2\theta$ from 0.001 to 0.5.
+At each grid point, the $\chi^{2}$ function is minimized with respect to three free parameters.
+One can then see how the goodness of fit varies over the grid, and how much of the grid is excluded based on a bad fit.
+By examining the excluded regions of the plane based on the high $\chi^{2}$ KATRIN concluded there was no evidence for sterile neutrinos at a high confidence level in their data over the entire grid.
+This negative result was surprising because the Neutrino-4 group had previously found evidence for a $2.70eV/c^{2}$ sterile neutrino in their neutrino oscillation experiment.
+Moreover, the KATRIN group is always very careful in their analysis of their data, so missing a signature of a sterile neutrino seems unlikely.
+Moreover, the conflict of KATRIN with the Neutrino-4 experiment was not a small one, since the data of the GALLEX, SAGE, and BEST experiments confirms the parameters of neutrino oscillations $\Delta m^{2}_{14}=7.3eV^{2}$ (or $m_{4}=2.70eV/c^{2}$ and $\sin^{2}2\theta_{14}=0.36)$ and the confidence level of the Neutrino-4 sterile neutrinos combined with the data of the GALLEX, SAGE, and BEST experiments is now $5.8\sigma.$ (Serebrov2024, )
+IV.10 Looking for the wrong kind of neutrinos?
+A likely possibility is that KATRIN failed to see any $m>1eV/c^{2}$ neutrinos because they were looking for the wrong kind of neutrinos in their sterile neutrino search.
+They should have been looking for active $m>1eV/c^{2}$ neutrinos not sterile ones – based on what was found in our SN 1987A data analysis.
+Any $m>1eV/c^{2}$ active neutrinos unfortunately would not show up given the way the KATRIN sterile search was conducted.
+Thus, obviously, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution.
+Rather they needed to fit their data using a combined differential spectrum $R_{C}$ for.the three masses $m_{1}=2.70eV/c^{2},m_{2}=22.4eV/c^{2}$ and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ with three weights, $\alpha_{j},$ that is:
+$R_{C}=\sum\alpha_{j}R_{\beta}(E,E_{0},m^{2}_{\nu,j})$ (6)
+where $R_{\beta}$ is the usual beta decay formula for a given $m^{2}$, E is the electron energy and $E_{0}$ is the fitted endpoint energy for tritium beta decay.
+For the weights $\alpha_{j}$ only two are independent since $\sum\alpha_{j}=1,$ and the $\alpha_{3}$ for the tachyonic mass is obviously very small in view of the very large magnitude of $|m_{3}^{2}|.$
+To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the spectrum are found – doing this both for the three stipulated masses and away from them.
+Finally, for the functional form of the spectrum for the tachyonic third mass, it may be possible to use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it.
+However, Ciborowski and Rembielinski have shown that for their tachyonic neutrino model a slightly more complicated formula is needed whose specific form depends on whether helicity or chirality coupling is used – see Eqs.
+20-21 in ref.
+(Ciborowski1999, )
+If KATRIN were to redo its analysis with these suggested changes it would be a decisive test of whether the SN 1989A data really can provide the three neutrino mass values, including the tachyonic one $m_{3}$.
+A positive KATRIN result which confirms all three masses from the SN 1987A data fit would be of some considerable interest, while a negative result would doom the idea of obtaining neutrino masses from the SN 1987A data.
+The missing $2.70eV/c^{2}$ sterile neutrinos in KATRIN was the key clue here, in the proverbial “dog that didn’t bark.” category.
+IV.11 Consistency with cosmological data?
+An apparent conflict exists between the three masses from the SN 1987A data fit and cosmological data.
+That conflict on the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ from cosmological data can be avoided, if one of the tachyon masses is negative gravitationally.
+(Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, meaning they have a repulsive gravitational effect.
+The current upper limit on the total sum of neutrino masses from cosmology has tightened to $\sum m_{\nu}<0.07-0.12eV/c^{2}$ (Shao2025, ), but some analyses of the data even favor a negative value for the sum of the neutrino masses.
+(Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found in the SN 1987A data, which could be avoided if the mass sum were over flavor state masses not mass states.
+In the recent literature the neutrino mass sum is considered to be the sum over mass state masses.
+However, going back to its original usage by Gershtein and Zeldovitch in 1966, the sum of the neutrino masses was intended to be a constraint on the sum of the flavor states – then only $\nu_{e}$ and $\nu_{\mu}.$ (Gershtein1966, ) However, at that time no distinction was made between mass and flavor states.
+Apart from historical usage there is a reason to view the sum as being over flavor state masses.
+Cosmological neutrinos were probably born in the universe’s infancy, decoupling from matter and radiation when the universe was only about 1 second old.
+At that time the two dominant processes capable of producing neutrinos was $Z^{0}$ decay or $e^{+}e^{-}$ collisions into neutrino antineutrino pairs, and would both produce equal numbers for each flavor.
+For example, $Z^{0}$ decays would result in equal $6.7\%$ branching ratios for each flavor, thus creating an equal number density N of each.
+Likewise for $e^{+}e^{-}$ collisions.
+In this case, the cosmological neutrino mass/energy density would be:
+$R=Nm_{\nu,e}+Nm_{\nu,\mu}+Nm_{\nu,\tau}=N\Sigma m_{\nu}$ (7)
+so, the appropriate constraint from Eq.
+7 should be on the flavor state sum, and not the mass state sum, as usually assumed.
+Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case.
+V Summary and the ultimate test
+This paper has made six claims that some readers may consider very difficult to believe – see Table 1.
+Some of the claims will likely not be accepted as true without the strongest corroborating evidence.
+After KATRIN does the analysis suggested in section IV.9, and it finds evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough.
 1. No valid basis exists to ignore the LSD 5-hr early burst
@@
 3. 3 $m_{j}^{2}$’s really can be found from 30 SN 1987A neutrinos
-4. Negative search results cannot rule out tachyon’s existence
 4. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
 5. The three $m^{2}=2.7^{2}eV^{2},22.4^{2}eV^{2}$ and $2\times 10^{5}eV^{2}$
-7. A likely explanation for these claims is CPT violation
-Table 1: The seven claims made in this paper are based on the SN 1987A neutrino data.
-Regarding number 4, not only are negative search results incapable of proving tachyonic neutrinos non-existent, but they do not even necessarily make them less likely, but they may narrow the range of possible negative $m^{2}$ values.
-We know SN 1987A was an asymmetric supernova (see Fig.5), with its expanding envelope and internal composition being lopsided rather than spherical.
-This spatial asymmetry was observed in the shape of the expanding envelope, and various other data.
-It is believed the asymmetry depicted likely arose from the progenitor star being a binary.
-Figure 6: The asymmetric distribution of the debris of an explosion of the core of a 23-solar-mass progenitor star at 4.3 seconds after core bounce.
-The coloring codes mass density, blue being higher density.
-Credit: Argonne National Lab
-Another particularly beautiful simulation of a supernova, this one lacking any symmetry whatsoever is shown in Fig.
-6.
-However, it is not a mere spatial asymmetry that might corroborate the claims made in this paper, but rather a much deeper symmetry violation: the breaking of CPT symmetry and the evidence we have found supporting that violation by a dimensionless amount $X_{CPT}=5\times 10^{-5}.$ Usually, violations of the fundamental symmetries involve a decisive experiment.
-Thus, for parity (P) this was the C.S.
-Wu experiment, (Wu1957, ) for CP symmetry, this was the Fitch-Cronin experiment.
-(Fitch1964, ) Should a future experiment demonstrate a breaking of CPT symmetry for neutrinos by an amount $X_{CPT}=5\times 10^{-5}$ that would certainly be very strong corroboration of the seven claims.
-However, to find truly unambiguous evidence for all the claims, there is only one (and likely only one) decisive test, namely a new galactic supernova.
+6. KATRIN likely can prove or reject all the above
+Table 1: six claims made in this paper.
+However, the most convincing test would be a new galactic supernova.
 Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors.
-The following concluding remarks represent one 88-year old’s supremely absurd optimism.
-When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other six claims.
-Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos, will definitely either confirm or reject the seven claims made.
-Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way to Earth from some star in our galaxy.
+This 88 year-old does not have a great chance of seeing it, but let’s be an optimist.
+When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other five claims.
+Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors.
+These neutrinos will definitely either confirm or reject the seven claims made.
+Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy.
 No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon.
-Soon in astronomical terms means potentially within the next 100,000 years (some say only 300 years).
-Were Betelgeuse to become a supernova, it would become incredibly bright, and even visible in daylight for many months.
-Go Betegeuse!
+Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )).
+Were Betelgeuse to become a supernova, it would become incredibly bright, and easily visible in daylight.
+Go Betelgeuse!
 Acknowledgments
 The author is grateful to Alan Chodos and Donald Gelman for helpful comments.
-The author would be incredibly grateful to any readers
-of this paper who find errors or poorly explained sections.
+If KATRIN performs the decisive test suggested in section IV.9 while I am still on this planet, I thank Guido Drexlin and Diana Parno of the KATRIN Group.
 Conflicts of interest

### v3

@@
 The key to finding the three neutrino masses is realizing why three normally accepted assumptions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any neutrino masses $m_{k}>1eV/c^{2}$ cannot be reconciled with upper limits on the “effective mass” from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A data is too great for the method to work.
-A particularly interesting piece of evidence supporting the claim made in the paper’s title involves a recent negative KATRIN result finding an absence of sterile neutrinos.
+A particularly crucial piece of evidence supporting the claim made in the paper’s title involves a recent negative KATRIN result finding an absence of sterile neutrinos.
 Keywords: SN 1987A, Neutrino, supernova, violation, tachyon, KATRIN, sterile neutrino
@@
 As a result, the author rediscovered what Ramanath Cowsik had found in 1988 about the neutrino masses being possibly revealed in the SN 1987A data.
-(Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) a clue from the KATRIN experiment’s reported failure to observe sterile neutrinos.
-This absence surprisingly strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation of the preposterous-sounding claim.
-(KATRIN2025, )
+(Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) an invaluable clue from the KATRIN experiment’s reported failure to observe sterile neutrinos.
+This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation of the preposterous-sounding claim.
+(KATRIN2025, ).
+The paper also suggests a way that short baseline $\nu_{e}$ disappearance experiments can supply such a confirmation.
 I.1 Are $m^{2}<0,$ and $v>c$ Tachyons “unphysical”?
@@
 Charles Schwartz is another theorist who has successfully challenged all the reasons why tachyons are considered unphysical, and shown how tachyons can account for many cosmological observations.
-(Schwartz2022, ; Schwartz2023, ).
+(Schwartz2018, ; Schwartz2022, ).
 I.2 Do negative searches make tachyons less likely?
@@
 Is it just a coincidence that the KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass value the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}?$ Ehrlich2015 Maybe it is a just a weird coincidence, but fortunately, as KATRIN takes more data the statistical uncertainty on the $m^{2}$ value will be reduced.
-Whatever the final KATRIN results, recall that the experiment measures the effective mass of the electron neutrino and not the three individual masses comprising it, which is the primary interest of this paper.
-Later we shall see how upper limits on the three neutrino masses from KATRIN and cosmological data can be circumvented.
+Whatever the final KATRIN results, recall that the experiment for now serves to set an upper limit on the effective mass of the electron neutrino and not the three individual masses comprising it, although that could change.
+Moreover, Later we shall see how the KATRIN upper limit can be circumvented.
 Without any specific a priori upper limits, it is conceivable the neutrino masses are large enough to be measurable based on the neutrino travel time over a very large distance.
@@
 (Agafonova2018, ) Nevertheless, even raising the possibility of tachyons probably was enough to dismiss the existence of the LSD neutrino burst as unrelated to SN 1987A in the minds of most physicists.
-On the other hand, there were four other reasons to be suspicious about the LSD burst besides the 5 hr early arrival, starting with it possibly being a statistical fluctuation of background.
+Moreover, there were four other reasons to be suspicious about the LSD burst besides the 5 hr early arrival, starting with it possibly being a statistical fluctuation of background.
 The five neutrinos observed during a 7 second time interval in LSD would, if not connected to SN1987A, represent a very significant departure from background (roughly one neutrino every 80 seconds), and that background had been steady for months.
@@
 Thus, the $\Omega^{-}$ particle which led to quarks was confirmed in 1964 at Brookhaven National Laboratory by observing just one single“Gold Standard” event.
-(Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the 30 neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only three points near the origin are unavoidably next to two lines.
+(Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the 30 neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only one data point very near the origin is within an error bar width of two lines.
 The least square fit of the 30 data points to three lines through the origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig.
@@
 Note the different scales used for $t>0$ and $t<0$ The five LSD data points at t=17ksect=-17ksec have been slightly separated from each other for clarity.
-The meaning of the shaded central area is discussed in the text
-The three masses corresponding to the slopes of the three straight lines in Fig.
+Based on Eq.
+5 the three masses corresponding to the slopes of the three straight lines in Fig.
 1 are $m_{1}=22.4eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ for the $m^{2}<0$ line.
-The author first discovered this three line fit in 2013, when he interpreted the result in terms of a 3 + 3 model (three pairs of active stable neutrinos – see Fig.3.) The big difference from the author’s previously published values of this fit (Ehrlich2013, ; Ehrlich2019, ) is in the value of $m_{2}$ because of the past omission of horizontal error bars.
+The author first discovered this three line fit in 2013, when he interpreted the result in terms of a 3 + 3 model (three pairs of active-sterile neutrinos – see Fig.3.) The big difference from the author’s previously published values of this fit (Ehrlich2013, ; Ehrlich2019, ) is in the value of $m_{2}$ because of the past omission of horizontal error bars.
 Those error bars are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line.
-We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.70eV/c^{2}$ and the $2.70eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration (both equal to within $1\%$), as well as its non-observation by KATRIN, (Serebrov2024, ; KATRIN2025, ) a matter discussed at length in section IV.9.
+We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.70eV/c^{2}$ and the $2.70eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration (both equal to within $1\%$), as well as its non-observation by KATRIN, (Serebrov2024, ; KATRIN2025, ) a matter discussed at length in section IV.10.
 III.1 Fixing a limitation of sterile neutrino searches
@@
 In effect, if you are only searching for one sterile neutrino when two or three are present they can mask each others presence.
-One way to avoid this problem is to search for two or three in the same fit to the data, and another way is to find a way to spatially separate any sterile neutrinos based on their masses, or find a place where nature causes this separation for us.
+One way to avoid this problem is to search for two or three in the same fit to the data, and another way is to find a way to spatially separate any sterile neutrinos based on their masses, or find a place where nature causes this separation for us – see next section.
 III.2 More evidence for the two $m^{2}>0$ masses
@@
 It seems unlikely one could find another case where nature is going to be so kind as to separate spatially different mass dark matter particles.
-However, as we shall see in section IV.9, KATRIN has a way to search for all three neutrino masses including the tachyonic one simultaneously in a single fit.
+However, as we shall see in section IV.10, KATRIN has a way to search for all three neutrino masses including the tachyonic one simultaneously in a single fit.
 Figure 2: The probability from the least squares fit to three straight lines through the origin versus the size of the horizontal error bars, assumed to be the same for all 30 data points.
 Note that there is no “best” fit to three straight lines unless one first fixes the size of the horizontal error bars because for $H>>5$ sec error bars the fit probability $p\approx 100\%$
-Figure 3: A possible $3+3$ neutrino mass squared diagram postulated by the author in 2013, showing updated $m_{1}$ and $m_{2}$ values.
+Figure 3: $3+3$ neutrino mass squared diagram postulated by the author in 2013, with updated $m_{1}$ and $m_{2}$ values.
 There are 3 active-sterile doublets (solid and dotted lines) split by three $\Delta m^{2}$ The three $\Delta m^{2}$ are for solar, atmospheric and short baseline (sbl) oscillations.
 Diagram is not to scale.
-The $m_{2}$ value differs significantly from the previously found value, i.e.
-$(m_{2}=4.0eV/c^{2})$ rather than $2.70eV/c^{2},$ because of the use of horizontal error bars now used in the three mass fit
+The “haunted” area is for $m^{2}<0$ neutrinos.
+Many physicists regard $m^{2}<0$ tachyons as ghost particles.
+Illustration of a ghostly nightmare is from Shutterstock.
 IV Discussion
@@
 This rejection was not because it provided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous.
-However since this issue and all the other reasons many people mistakenly have for rejecting the LSD data can be answered, there is no valid reason not to keep LSD data, especially because with three neutrinos the mandated number of straight lines about which clustering should occur is three.
+However since this issue and all the other reasons many people mistakenly have for rejecting the LSD data can be answered, there is no valid reason not to keep LSD data, especially because with three active neutrinos the mandated number of straight lines about which clustering should occur is three.
 IV.2 Why are the LSD data so bunched in time?
@@
 IV.4 Alternate explanations for the LSD neutrinos?
-Many phenomena in nature on the fringes of the known have alternative explanations. All the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems. Thus,
-  • If it was a statistical fluctuation it was a really big one $p=1.4\times 10^{-6}$
-  • If it was due to a double bang the pattern in Fig. 1 should look very different as discussed in section IV.7.
-  • If it was a genuine, low-energy pulse (e.g., from neutron capture on iron slabs around the detector), it would require the supernova to release an unusually high amount of energy in this form.
+Many phenomena in nature on the fringes of the known have alternative explanations.
+All the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems.
+Thus, perhaps the most likely alternate possibility is that a supernova can bang more than once.
+In fact, Galeotti and Pizzella have even claimed there was a third bang– not in the LSD detector but in Kamiokande II at a time about 20 min after the main burst.
+(Galeotti2007, ) However, Kamiokande II does not acknowledge this burst.
+Supernovae undergoing multiple bangs is a known phenomenon (known as Pulsational pair-instability supernovae), which can occur for very massive stars according to Woosley.
+(Woosley2017, ) According to him SN 1987A had a helium core of 5 to 6 $M_{sun}$ which is characteristic of an $18-20M_{sun}$ main sequence star.
+However, this mass is too small to experience the pair instability which requires He core masses of at least $37M_{sun}.$ Additionally, Woosley notes that pulsational pair instaility supernovae do not produce detectable neutrino bursts.
+Among the other alternative explanations of the LSD burst are:
+  • If it was a statistical fluctuation it was a really big one $p=1.4\times 10^{-6}.$ In fact among the LSD signal candidates for the detection of a neutrino burst selected over 14 years of operation, it has the lowest background imitation probability.
+  • If it was a genuine, low-energy pulse (e.g., from neutron capture on iron slabs around the detector), simulations show that their number is completely negligible for all of the neutrino radiation parameters considered.
+(Manukovskiy2022, ).
   • If it was due to 8 MeV dark matter as suggested in ref.
 (Ehrlich2021, ) this would still create the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD neutrinos falling on the $m^{2}<0$ line.
+  • The only other explanation known to this author for the LSD burst besides tachyons is the Frogatt-Nielsen proposal of “dark matter balls” inside the the stellar core (Froggatt2015, ).
 IV.5 Best size of the horizontal error bars?
@@
 (Bozza2025, ).
-Note the time scale change at $t=1$ sec.
-The fraction of the integrated luminosity for $t<1$ is less than it seems because of the scale change.
+Note the time scale change at $t=1$ seconds.
+The fraction of the integrated luminosity for $t<1$ seconds is only a bit more than half because of that scale change.
 Figure is used under creative commons license https://creativecommons.org/licenses/by/4.0
-IV.7 Suppose neutrinos all have $m<<1eV/c^{2}$?
+IV.7 Suppose neutrinos have $m<<1eV/c^{2}$?
 It is natural to wonder what Fig.
-1 would look like if the three neutrino masses were very small and none were tachyonic, as in the standard neutrino description.
-Clearly, they still would lie on three lines, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics.
-However, we almost certainly would not be able to find these lines because for 3 small masses we would find three nearby steeply sloped lines, as in Fig.5.
-showing how the 3 lines might appear for neutrinos emitted at $t=0$ and $t=1$ sec Of course, with a horizontal error bar of 1 or 2 sec, identifying which of 30 points went with which nearly vertical line would be impossible because there are in fact many more line-triplets between $t=0$ and $t=1sec.$
-If the neutrino mass were in fact under the KATRIN current upper limit $(0.45eV/c^{)}$ the data points would all lie on lines that were virtually vertical and the data points would all lie in the shaded central area in Fig.1 having a 2 second time window.
-This is obviously far from reality, which suggests that their mass is not below the KATRIN limit.
-In the “double bang” scenario for SN 1987A which could conceivably explain the LSD burst, the preceding remains true, so once again we would not be able to identify the three lines about which the 30 points are clustering, unlike the actual data in Fig.
-1.
-This means that the double bang scenario can be ruled out, given Fig.1.
-Figure 5: The case of neutrinos having very small masses (which yield triplets of steeply sloped straight lines nearly overlapping) with a data point having a 2 sec wide horizontal error bar.
-Neutrinos are assumed emitted at $t=0$ and $t=1$ seconds.
-Given the size of the error bar, it would be impossible to find which of the six near vertical lines the data point is associated with (likewise for all the other line triplets for $0>t>1$ seconds – completely unlike what is actually the case in Fig.1.
-IV.8 Consistency with direct mass experiments?
+1 would look like if the three neutrino masses were small and none were tachyonic, as in the standard neutrino description.
+Clearly, they still would lie on three lines if emitted simultaneously, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics.
+However, we would not be able to find these lines because for 3 small masses they would lie on very many almost vertical lines filling the time window – see the 2 second error bar in Fig, 5.
+In effect each point lies on its own vertical line filling the time window.
+This is obviously far from what is the case for the actual data, which suggests that the neutrino mass does not satusfy $m<<1eV/c^{2}.$
+Figure 5: How 50 neutrinos might appear on a $1/E^{2}$ versus t plot if their mass $m<<1eV/c^{2}$ and they were emitted during a time window of 2 seconds.
+No congregation of the points about specific lines would be observed, completely unlike the real SN 1987A data which clearly clusters along three straight lines.
+Essentially, each of the 50 points lies on its own different vertical line somewhere within the 2 second time window and no two points lie on the same line.
+IV.8 Consistency with cosmological data?
+An apparent conflict exists between the three masses from the SN 1987A data fit and cosmological data.
+That conflict on the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ from cosmological data can be avoided, if one of the tachyon masses is negative gravitationally.
+(Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, meaning they have a repulsive gravitational effect.
+The current upper limit on the total sum of neutrino masses from cosmology has tightened to $\sum m_{\nu}<0.07-0.12eV/c^{2}$ (Shao2025, ), but some analyses of the data even favor a negative value for the sum of the neutrino masses.
+(Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found in the SN 1987A data, which could be avoided if the mass sum were over flavor state masses not mass states.
+In the recent literature the issue of whether the sum is over mass or flavor states is not widely discussed.
+However, going back to its original usage by Gershtein and Zeldovitch in 1966, the sum of the neutrino masses was intended to be a constraint on the sum of the flavor states – then only $\nu_{e}$ and $\nu_{\mu}.$ (Gershtein1966, ) However, at that time no distinction was made between mass and flavor states.
+Apart from historical usage there is a reason to view the sum as being over flavor state masses.
+Cosmological neutrinos were probably born in the universe’s infancy, decoupling from matter and radiation when the universe was only about 1 second old.
+At that time the two dominant processes capable of producing neutrinos was $Z^{0}$ decay or $e^{+}e^{-}$ collisions into neutrino antineutrino pairs, and would both produce equal numbers for each flavor.
+For example, $Z^{0}$ decays would result in equal $6.7\%$ branching ratios for each flavor, thus creating an equal number density N of each.
+Likewise for $e^{+}e^{-}$ collisions.
+Thus, the cosmological neutrino mass/energy density produced in that manner would be:
+$R=Nm_{\nu,e}+Nm_{\nu,\mu}+Nm_{\nu,\tau}=N\Sigma m_{\nu}$ (6)
+so, the appropriate constraint on R from fitting cosmological data is from Eq.
+6 a constraint on the flavor state sum, and not the mass state sum, as sometimes assumed.
+Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case.
+IV.9 Consistency with direct mass experiments?
 The three masses deduced from the three straight line fit in Fig.
-1 seem preposterously large, and they would appear to be inconsistent with data, whether it be cosmological data or data from the KATRIN experiment implying $m_{\nu}<0.45eV/c^{2}$.
+1 seem preposterously large, and they would appear to be inconsistent with data from the KATRIN experiment implying $m_{\nu}<0.45eV/c^{2}$.
 As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the electron neutrino effective mass has $m_{k}^{2}<0$ the effective mass, as normally defined, $m_{e}^{2}=\sum|U_{e,k}|^{2}m_{k}^{2}$ can be arbitrarily close to zero or even negative, even when the $m_{k}^{2}$ are all very large, as in the SN 1987 fit.
-KATRIN has further crucial relevance to that fit quite apart from the limit it imposes by its measured $\nu_{e}$ effective mass.
-IV.9 A missing $m_{4}=2.70eV/c^{2}$ sterile neutrino?
+KATRIN has further crucial relevance quite apart from the limit it imposes by its measured $\nu_{e}$ effective mass.
+IV.10 A missing $m_{4}=2.70eV/c^{2}$ sterile neutrino?
 KATRIN recently reported finding no evidence for sterile neutrinos in its data based on the observed shape of the tritium beta decay spectrum.
@@
 Moreover, the conflict of KATRIN with the Neutrino-4 experiment was not a small one, since the data of the GALLEX, SAGE, and BEST experiments confirms the parameters of neutrino oscillations $\Delta m^{2}_{14}=7.3eV^{2}$ (or $m_{4}=2.70eV/c^{2}$ and $\sin^{2}2\theta_{14}=0.36)$ and the confidence level of the Neutrino-4 sterile neutrinos combined with the data of the GALLEX, SAGE, and BEST experiments is now $5.8\sigma.$ (Serebrov2024, )
-IV.10 Looking for the wrong kind of neutrinos?
-A likely possibility is that KATRIN failed to see any $m>1eV/c^{2}$ neutrinos because they were looking for the wrong kind of neutrinos in their sterile neutrino search.
-They should have been looking for active $m>1eV/c^{2}$ neutrinos not sterile ones – based on what was found in our SN 1987A data analysis.
-Any $m>1eV/c^{2}$ active neutrinos unfortunately would not show up given the way the KATRIN sterile search was conducted.
-Thus, obviously, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution.
-Rather they needed to fit their data using a combined differential spectrum $R_{C}$ for.the three masses $m_{1}=2.70eV/c^{2},m_{2}=22.4eV/c^{2}$ and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ with three weights, $\alpha_{j},$ that is:
+IV.11 “The dog that didn’t bark” (the key section!)
+It is possible that KATRIN failed to see any sterile neutrinos because they simply do not exist, despite the above high confidence kevel, but here we consider a much more interesting possibility.
+Neutrino - 4 and the other three experiments that saw the $2.70eV/c^{2}$ neutrino were doing a short baseline disappearance oscillation experiment, which observed the disappearance probability $P(\bar{\nu_{e}}\rightarrow\bar{\nu_{e}}).$ This type of experiment would work just as well if the disappearing neutrinos were disappearing into sterile or active neutrinos.
+However, KATRIN’s search was specifically looking for sterile neutrinos and it would not work for active ones as conducted.
+(KATRIN was not being careless here, they just assumed like everyone else that neutrinos with large mass would have to be sterile) Of course, if the main claim of this paper is correct, and if the $2.70eV/c^{2}$ neutrino is an active neutrino (as well as a sterile one - see Fig.
+4), then KATRIN’s failure to see it is understandable because their search method just would not work for active neutrinos.
+Specifically, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution.
+(Why not you ask?
+Because the zero mass effective mass is made up of the three specific masses we found including one with $m^{2}<0$) Rather, they needed to fit their data using a combined differential spectrum $R_{C}$ for.the three masses $m_{1}=2.70eV/c^{2},m_{2}=22.4eV/c^{2}$ and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ with three unknown adjustible weights, $\alpha_{j},$ that is:
 $R_{C}=\sum\alpha_{j}R_{\beta}(E,E_{0},m^{2}_{\nu,j})$ (7)
@@
 For the weights $\alpha_{j}$ only two are independent since $\sum\alpha_{j}=1,$ and the $\alpha_{3}$ for the tachyonic mass is obviously very small in view of the very large magnitude of $|m_{3}^{2}|.$
-To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the spectrum are found – doing this both for the three stipulated masses and away from them.
-Finally, for the functional form of the spectrum for the tachyonic third mass, it may be possible to use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it.
-However, Ciborowski and Rembielinski have shown that for their tachyonic neutrino model a slightly more complicated formula is needed whose specific form depends on whether helicity or chirality coupling is used – see Eqs.
-20-21 in ref.
-(Ciborowski1999, )
-If KATRIN were to redo its analysis with these suggested changes it would be a decisive test of whether the SN 1989A data really can provide the three neutrino mass values, including the tachyonic one $m_{3}$.
-A positive KATRIN result which confirms all three masses from the SN 1987A data fit would be of some considerable interest, while a negative result would doom the idea of obtaining neutrino masses from the SN 1987A data.
-The missing $2.70eV/c^{2}$ sterile neutrinos in KATRIN was the key clue here, in the proverbial “dog that didn’t bark.” category.
-IV.11 Consistency with cosmological data?
-An apparent conflict exists between the three masses from the SN 1987A data fit and cosmological data.
-That conflict on the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ from cosmological data can be avoided, if one of the tachyon masses is negative gravitationally.
-(Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, meaning they have a repulsive gravitational effect.
-The current upper limit on the total sum of neutrino masses from cosmology has tightened to $\sum m_{\nu}<0.07-0.12eV/c^{2}$ (Shao2025, ), but some analyses of the data even favor a negative value for the sum of the neutrino masses.
-(Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found in the SN 1987A data, which could be avoided if the mass sum were over flavor state masses not mass states.
-In the recent literature the neutrino mass sum is considered to be the sum over mass state masses.
-However, going back to its original usage by Gershtein and Zeldovitch in 1966, the sum of the neutrino masses was intended to be a constraint on the sum of the flavor states – then only $\nu_{e}$ and $\nu_{\mu}.$ (Gershtein1966, ) However, at that time no distinction was made between mass and flavor states.
-Apart from historical usage there is a reason to view the sum as being over flavor state masses.
-Cosmological neutrinos were probably born in the universe’s infancy, decoupling from matter and radiation when the universe was only about 1 second old.
-At that time the two dominant processes capable of producing neutrinos was $Z^{0}$ decay or $e^{+}e^{-}$ collisions into neutrino antineutrino pairs, and would both produce equal numbers for each flavor.
-For example, $Z^{0}$ decays would result in equal $6.7\%$ branching ratios for each flavor, thus creating an equal number density N of each.
-Likewise for $e^{+}e^{-}$ collisions.
-In this case, the cosmological neutrino mass/energy density would be:
-$R=Nm_{\nu,e}+Nm_{\nu,\mu}+Nm_{\nu,\tau}=N\Sigma m_{\nu}$ (7)
-so, the appropriate constraint from Eq.
-7 should be on the flavor state sum, and not the mass state sum, as usually assumed.
-Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case.
+To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the observed spectrum are found.
+With the three masses set to their specified SN 1987A values, KATRIN should observe a small concentrated area where good fits are found surrounded by a large area of poor fits.
+They could also do the same using other masses, where no good fits should be found.
+The best fit on the $\alpha_{1}$ and $\alpha_{2}$ plane should yield the effective mass from $m^{2}=\sum\alpha_{j}m^{2}_{j},$ and it should be a very small negative value.
+Finally, for the functional form of the spectrum for the tachyonic third mass, $m_{3}^{2},$ they could just use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it.
+While Rembielinski, Caban and Ciborowski have shown that their causal tachyonic neutrino model results in a more complicated formula – see Eqs.
+93-94 in ref.
+(Rembielinski2021, ), however, as shown in their Fig.
+1 the simple approach gives a result that is virtually indistinguishable from their full description.
+In summary, if KATRIN were to redo its analysis with the preceding suggested changes it could be a decisive test of whether the SN 1989A data really can provide the three neutrino mass values, including the tachyonic one $m_{3}$.
+A positive KATRIN result which confirms all three masses from the SN 1987A data fit would be of some considerable interest.
+The missing $2.70eV/c^{2}$ sterile neutrinos in the KATRIN search was the key clue here.
+It was effectively “the dog that didn’t bark,” a crucial example of a “negative fact.”
+IV.12 Finding more evidence in oscillation data?
+Neutrino oscillation experiments of the dissapearing type can also be used to find evidence of the three neutrino masses from the SN 1987A data using a similar approach.
+The key quantity fitted to the data in such an experiment is the survival probability $P_{j}$ of a neutrino of a given flavor.
+One usually looks to see if there is any significant amplitude $\alpha_{j}$ for the probability anywhere on a grid of values over sterile neutrino mass and amplitude of the oscillation for a single sterile neutrino.
+To find how much amplitude oscillation there is in the data for neutrinos of mass $m_{1}=2.70eV/c^{2}$ and simultaneously for neutrinos of mass $m_{2}=22.4eV/c^{2}$ one can create a grid of $\alpha_{1}$ and $\alpha_{2}$ values with the summed probability $P_{1}+P_{2}$ calculated at each grid point, and see if there is a small region of the grid where the summed probabilities is very high.
+If so, that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data.
+The most promising experiments to look at might be the four identified in ref.
+(Serebrov2024, ) which already have shown evidence of the $2.70eV/c^{2}$ mass – do they now show evidence of a $22.4eV/c^{2}$ neutrino as well?
+Of course, only some experiments having a very short baseline will be sensitive to a mass this high.
+It would also be interesting to look at some formerly negative results such as one from the NOvA Collaboration to see if they see anything.
+(Acero2025, )
 V Summary and the ultimate test
-This paper has made six claims that some readers may consider very difficult to believe – see Table 1.
+This paper has made six claims that many.
+readers may understandably consider very difficult to believe – see Table 1.
 Some of the claims will likely not be accepted as true without the strongest corroborating evidence.
-After KATRIN does the analysis suggested in section IV.9, and it finds evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough.
+After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough.
+This might also be the case if dissapearance experiments were to find such evidence doing the analysis mentioned in the previous section.
 1. No valid basis exists to ignore the LSD 5-hr early burst
@@
 5. The three $m^{2}=2.7^{2}eV^{2},22.4^{2}eV^{2}$ and $2\times 10^{5}eV^{2}$
-6. KATRIN likely can prove or reject all the above
+6. Existing data likely can prove or reject all the above
 Table 1: six claims made in this paper.
@@
 Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors.
-These neutrinos will definitely either confirm or reject the seven claims made.
+These neutrinos will definitely either confirm or reject the six claims made.
 Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy.
@@
 Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )).
-Were Betelgeuse to become a supernova, it would become incredibly bright, and easily visible in daylight.
+Were Betelgeuse to become a supernova, it would become incredibly bright, potentially as bright as the full moon.
 Go Betelgeuse!
 Acknowledgments
-The author is grateful to Alan Chodos and Donald Gelman for helpful comments.
-If KATRIN performs the decisive test suggested in section IV.9 while I am still on this planet, I thank Guido Drexlin and Diana Parno of the KATRIN Group.
+The author is grateful to Alan Chodos and Donald Gelman for helpful comments. I also thank Guido Drexlin and Diana Parno of the KATRIN Collaboration for their kindness.
 Conflicts of interest

### v4

@@
 (Ehrlich2021, ) this would still create the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD neutrinos falling on the $m^{2}<0$ line.
-  • The only other explanation known to this author for the LSD burst besides tachyons is the Frogatt-Nielsen proposal of “dark matter balls” inside the the stellar core (Froggatt2015, ).
+  • The only other explanation known to this author for the LSD burst besides tachyons is the Froggatt-Nielsen proposal of “dark matter balls” inside the the stellar core (Froggatt2015, ).
 IV.5 Best size of the horizontal error bars?
@@
 It was effectively “the dog that didn’t bark,” a crucial example of a “negative fact.”
-IV.12 Finding more evidence in oscillation data?
-Neutrino oscillation experiments of the dissapearing type can also be used to find evidence of the three neutrino masses from the SN 1987A data using a similar approach.
-The key quantity fitted to the data in such an experiment is the survival probability $P_{j}$ of a neutrino of a given flavor.
-One usually looks to see if there is any significant amplitude $\alpha_{j}$ for the probability anywhere on a grid of values over sterile neutrino mass and amplitude of the oscillation for a single sterile neutrino.
-To find how much amplitude oscillation there is in the data for neutrinos of mass $m_{1}=2.70eV/c^{2}$ and simultaneously for neutrinos of mass $m_{2}=22.4eV/c^{2}$ one can create a grid of $\alpha_{1}$ and $\alpha_{2}$ values with the summed probability $P_{1}+P_{2}$ calculated at each grid point, and see if there is a small region of the grid where the summed probabilities is very high.
-If so, that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data.
-The most promising experiments to look at might be the four identified in ref.
-(Serebrov2024, ) which already have shown evidence of the $2.70eV/c^{2}$ mass – do they now show evidence of a $22.4eV/c^{2}$ neutrino as well?
-Of course, only some experiments having a very short baseline will be sensitive to a mass this high.
-It would also be interesting to look at some formerly negative results such as one from the NOvA Collaboration to see if they see anything.
-(Acero2025, )
+IV.12 Multiple sterile neutrinos in oscillation data?
+Anatolii Serebrov and his Neutrino 4 team, who first detected reactor antineutrino oscillations in the 6-12 meter range, has advocated the need for two more sterile neutrinos in addition to $m_{1}^{2}=2.70^{2}eV^{2}/c^{4}$ based on theoretical grounds (Serebrov2024, ) Assuming a second sterile neutrino of mass $m_{2}^{2}=22.4eV^{2}/c^{2},$ the oscillations would presumably show up in a range roughly 80 times shorter.
+If such oscillations are detectable one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find thelikelihood probability of the fit to the data for that pair of mass values.
+One could then see where on the plane that likelihood maximizes.
+If there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},22.4^{2})eV^{2}/c^{2},$ that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data.
+The most promising experiments to look at might be the four identified by Serebrov in ref.
+(Serebrov2024, ).
+These already have shown evidence of the $2.70eV/c^{2}$ mass, and the question is do they now also show evidence of a $22.4eV/c^{2}$ neutrino.
+It would seem highly unlikely that a signature of any oscillation due to the $m_{3}^{2}=-200,000eV^{2}/c^{4}$ neutrino could be found in a similar manner, but it should be detectable in KATRIN as already discussed.
 V Summary and the ultimate test
@@
 After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough.
-This might also be the case if dissapearance experiments were to find such evidence doing the analysis mentioned in the previous section.
-1. No valid basis exists to ignore the LSD 5-hr early burst
-2. No a priori upper limits exist on the three neutrino $m^{2}_{\nu,j}$
+This might also be the case if oscillation experiments were to find such evidence doing the analysis mentioned in the previous section.
+1. The LSD 5-hr early burst was connected to SN 1987A
+2. Upper limits on the 3 neutrino $m^{2}_{\nu,j}$ can be avoided
 3. 3 $m_{j}^{2}$’s really can be found from 30 SN 1987A neutrinos
@@
 5. The three $m^{2}=2.7^{2}eV^{2},22.4^{2}eV^{2}$ and $2\times 10^{5}eV^{2}$
-6. Existing data likely can prove or reject all the above
+6. Existing data likely can confirm or reject all the above
 Table 1: six claims made in this paper.

### v5

@@
 A particularly crucial piece of evidence supporting the claim made in the paper’s title involves a recent negative KATRIN result finding an absence of sterile neutrinos.
-Keywords: SN 1987A, Neutrino, supernova, violation, tachyon, KATRIN, sterile neutrino
+This absence of a sterile neutrino signal leads to two tests of the claim based on existing data: one for KATRIN and one for IceCube.
+Keywords: SN 1987A, Neutrino, supernova, tachyon, KATRIN, IceCube, sterile neutrino
 I Introduction
@@
 (KATRIN2025, ).
-The paper also suggests a way that short baseline $\nu_{e}$ disappearance experiments can supply such a confirmation.
+The paper also suggests a way that oscillation experiments and especially the IceCube experiment might supply such a confirmation.
 I.1 Are $m^{2}<0,$ and $v>c$ Tachyons “unphysical”?
@@
 Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case.
+Conversely, for the neutrino mass sum to be a sum over mass states, there would need to be a process that created neutrinos in equal numbers according to their mass state, and there is no such process.
 IV.9 Consistency with direct mass experiments?
@@
 To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the observed spectrum are found.
-With the three masses set to their specified SN 1987A values, KATRIN should observe a small concentrated area where good fits are found surrounded by a large area of poor fits.
+With the three masses set to their specified SN 1987A values, KATRIN should ideally observe a small concentrated area where good fits are found surrounded by a large area of poor fits.
 They could also do the same using other masses, where no good fits should be found.
 The best fit on the $\alpha_{1}$ and $\alpha_{2}$ plane should yield the effective mass from $m^{2}=\sum\alpha_{j}m^{2}_{j},$ and it should be a very small negative value.
-Finally, for the functional form of the spectrum for the tachyonic third mass, $m_{3}^{2},$ they could just use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it.
+Finally, for the functional form of the spectrum for the tachyonic third mass, $m_{3}^{2},$ KATRIN could just use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it.
 While Rembielinski, Caban and Ciborowski have shown that their causal tachyonic neutrino model results in a more complicated formula – see Eqs.
 93-94 in ref.
-(Rembielinski2021, ), however, as shown in their Fig.
+(Rembielinski2021, ), as shown in their Fig.
 1 the simple approach gives a result that is virtually indistinguishable from their full description.
@@
 IV.12 Multiple sterile neutrinos in oscillation data?
-Anatolii Serebrov and his Neutrino 4 team, who first detected reactor antineutrino oscillations in the 6-12 meter range, has advocated the need for two more sterile neutrinos in addition to $m_{1}^{2}=2.70^{2}eV^{2}/c^{4}$ based on theoretical grounds (Serebrov2024, ) Assuming a second sterile neutrino of mass $m_{2}^{2}=22.4eV^{2}/c^{2},$ the oscillations would presumably show up in a range roughly 80 times shorter.
-If such oscillations are detectable one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find thelikelihood probability of the fit to the data for that pair of mass values.
+Anatolii Serebrov and his Neutrino 4 team, have argued on cosmological grounds for there being two heavier sterile neutrinos in addition to the one found at $m_{1}^{2}=2.70^{2}eV^{2}/c^{4},$ which of course was also earlier inferred from the SN 1987A data analysis.
+Assuming a second sterile neutrino of mass $m_{2}^{2}=22.4eV^{2}/c^{2},$ as mandated by that data, the oscillation wavelength would be 69 times shorter than the $m_{1}$ sterile neutrinos.
+It is highly unlikely that such short wavelength oscillations are detectable in the data of Neutrino 4, but if they were one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find the likelihood probability of the fit to the data for that pair of mass values.
 One could then see where on the plane that likelihood maximizes.
 If there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},22.4^{2})eV^{2}/c^{2},$ that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data.
-The most promising experiments to look at might be the four identified by Serebrov in ref.
-(Serebrov2024, ).
-These already have shown evidence of the $2.70eV/c^{2}$ mass, and the question is do they now also show evidence of a $22.4eV/c^{2}$ neutrino.
-It would seem highly unlikely that a signature of any oscillation due to the $m_{3}^{2}=-200,000eV^{2}/c^{4}$ neutrino could be found in a similar manner, but it should be detectable in KATRIN as already discussed.
+You might think that the most promising experiments to look at would be the four data sets identified by Serebrov in ref.
+(Serebrov2024, ), which already have shown evidence of the $2.70eV/c^{2}$ mass at a cumulative $5.8\sigma$.
+Unfortunately, however, the very short wavelength of $22.4^{2}eV^{2}/c^{2}$ neutrinos almost certainly make these oscillations undetectable in existing short baseline reactor data.
+There may be over 20 projects in the works exploring the ultra-short baseline region ($L<10$ meters) in reactor experiments, (Serebrov2021, ) but no experiments known to the author could detect stable sterile neutrinos of such a high mass, with the exception of IceCube.
+In fact, IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in seven years of data.
+(Abbasi2024, ) Even though they reported a negative result their best fit did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass whuch was strongly preferred over the no oscillation solution by 10.7 to one.
+Furthermore, the IceCube energy range (up to $E=100TeV$) would make them sensitive to sterile mass values as high as $m_{2}^{2}=22.4^{2}eV^{2}/c^{4}.$ Thus, it might make sense for them to rexamine their data for evidence of a double oscillation for those two specific masses.
+Making such a search even more promissing, IceCube has already begun to explore the region of $\Delta m^{2}>10eV^{2}/c^{4},$ and found a preference for $3+1$ oscillations.
+(Biennow2018, ) A search using the method outlined for the Neutrino-4 data might prove particularly promising.
+Although the SN 1987A data analysis specified three neutrino masses it seems highly unlikely that the $m_{3}^{2}=-200,000eV^{2}/c^{4}$ neutrino would leave any detectable signal in IceCube or any other oscillation data (unlike KATRIN where it should be detectable).
 V Summary and the ultimate test
-This paper has made six claims that many.
-readers may understandably consider very difficult to believe – see Table 1.
+This paper has made six claims that many readers may understandably consider very difficult to believe – see Table 1.
 Some of the claims will likely not be accepted as true without the strongest corroborating evidence.
-After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough.
+After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data, that might be enough.
 This might also be the case if oscillation experiments were to find such evidence doing the analysis mentioned in the previous section.
-1. The LSD 5-hr early burst was connected to SN 1987A
+1. The LSD burst was due to SN 1987A but no 2nd bang
 2. Upper limits on the 3 neutrino $m^{2}_{\nu,j}$ can be avoided
@@
 5. The three $m^{2}=2.7^{2}eV^{2},22.4^{2}eV^{2}$ and $2\times 10^{5}eV^{2}$
-6. Existing data likely can confirm or reject all the above
+6. KATRIN, IceCube or other data likely can confirm #5
 Table 1: six claims made in this paper.

### v6

@@
 A particularly crucial piece of evidence supporting the claim made in the paper’s title involves a recent negative KATRIN result finding an absence of sterile neutrinos.
-This absence of a sterile neutrino signal leads to two tests of the claim based on existing data: one for KATRIN and one for IceCube.
+This absence of a sterile neutrino signal leads to two tests of the claim based on existing data in KATRIN and IceCube.
 Keywords: SN 1987A, Neutrino, supernova, tachyon, KATRIN, IceCube, sterile neutrino
@@
 About $10^{58}$ neutrinos were emitted from the dying star Sanduleak 69 202 when it became the first supernova seen in 1987 (hence the name SN 1987A).
-Of those $10^{58}$ neutrinos only a couple of dozen were captured by Earth’s detectors.
-Yet, thousands of papers have been published about SN 1987A, and its well-studied nature is not surprising given that it is the only supernova in our galaxy since neutrino detectors first existed.
+Of those $10^{58}$ neutrinos only a few dozen were captured by Earth’s detectors.
+Yet, thousands of papers have been published about SN 1987A, and its well-studied nature is not surprising given that it is the only supernova in or close to our galaxy since neutrino detectors first existed.
 The earliest articles on the neutrino mass obtained from SN 1987A data, with one exception,(Cowsik1988, ) set only upper limits on its value.
@@
 (Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) an invaluable clue from the KATRIN experiment’s reported failure to observe sterile neutrinos.
-This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation of the preposterous-sounding claim.
 (KATRIN2025, ).
+This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation or rejection of the preposterous-sounding claim suggested in the title.
 The paper also suggests a way that oscillation experiments and especially the IceCube experiment might supply such a confirmation.
@@
 I.2 Do negative searches make tachyons less likely?
-The question of whether $m^{2}<0$ and $v>c$ observable particles exist is not theoretical, but strictly empirical.
-All reports of tachyons, including the 2011 OPERA experiment initial $v>c$ result, (Brumfiel2012, ) have later been shown to be incorrect apart from one-time sightings.
-(Ehrlich2022, ) Most physicists therefore believe that the tachyon hypothesis can be dismissed, or at least relegated to the highly implausible category.
+Most reports of tachyons, including the 2011 OPERA experiment initial $v>c$ result, (Brumfiel2012, ) have later been shown to be incorrect apart from one-time sightings.
+(Ehrlich2022, ) Many physicists therefore believe that the tachyon hypothesis can be dismissed, or at least relegated to the highly implausible category.
 A more reasonable conclusion is that as long as not a single neutrino has ever been measured to have either $m^{2}>0$ or $v<c,$ the tachyonic neutrino hypothesis remains viable.
 A decline in belief in physical tachyons is perfectly understandable as a psychological reaction to negative results or even worse, false positive results like OPERA.
-However, negative and false positive results do not suggest tachyons do not exist, only that the $m^{2}$ for neutrinos is so close to zero, that we cannot tell if $m^{2}>0$ or $m^{2}<0$.
+However, negative and false positive results do not suggest tachyons do not exist, only that the $m^{2}$ for neutrinos is so close to zero, that we cannot yet tell if $m^{2}>0$ or $m^{2}<0$.
 You might think a negative hypothetical result $m^{2}=0.0\pm 0.1eV^{2}/c^{4}$ makes tachyonic neutrinos more unlikely than a previous result $m^{2}=0.0\pm 10.0eV^{2}/c^{4}$ because the allowed negative range in $m^{2}$ has been reduced a hundredfold.
@@
 Neutrinos of course are the only known particles that might be tachyons since all others have been observed travelling at speeds $v<c,$ and tachyons if they exist must always be superluminal.
-I.3 KATRIN and circumventing upper limits on $m_{\nu}.$
-The KATRIN experiment seeks to find the “effective mass” of the electron antineutrino $m^{2}=\Sigma|U_{i,j}|^{2}m_{j}^{2}.$ Its best value as of 2025 is negative: $m^{2}=-0.14^{+0.13}_{-0.15}eV^{2}.$ (KATRIN2024, ) This result, of course, leaves the true sign of $m^{2}$ ambiguous given the size of the uncertainty.
-Is it just a coincidence that the KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass value the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}?$ Ehrlich2015 Maybe it is a just a weird coincidence, but fortunately, as KATRIN takes more data the statistical uncertainty on the $m^{2}$ value will be reduced.
-Whatever the final KATRIN results, recall that the experiment for now serves to set an upper limit on the effective mass of the electron neutrino and not the three individual masses comprising it, although that could change.
-Moreover, Later we shall see how the KATRIN upper limit can be circumvented.
-Without any specific a priori upper limits, it is conceivable the neutrino masses are large enough to be measurable based on the neutrino travel time over a very large distance.
-The best baseline to measure the three masses based on their travel time is almost certainly a galactic supernova, provided we are confident the neutrino emissions from it are nearly simultaneous.
-An extragalactic one might seem even better were it not for the inverse square law and the very few neutrinos that could reach us.
-The phrase “neutrino burst” is used throughout this paper, referring to a sudden, intense emission of a vast number of neutrinos from a powerful cosmic event, most notably a core-collapse supernova lasting seconds.
-However we also use the phrase to refer to a seconds-long grouping of many more neutrinos than normally seen in a neutrino detector in such a time interval, which are believed to have originated from a supernova.
+I.3 Assumptions in neutrino mass experiments
+To date most experiments to measure the “effective mass” of the electron neutrino (or antineutrino) have done so by fitting the shape of the beta decay spectrum of tritium ${{}^{3}H}\rightarrow{{}^{3}He}+e+\bar{\nu_{e}}$ very close to its endpoint.
+It is assumed that the “effective mass” of the electron neutrino is defined in terms of the three states having masses $m_{j}$ and the PMNS mixing matrix $U_{i,j}$ as
+$m_{\nu}^{2}=\Sigma|U_{i,j}|^{2}m_{j}^{2}$ (1)
+For the electron, muon and tau neutrinos $i=1,2,3,$ respectively in Eq.
+1.
+Many experiments in the past have reported negative values for $m_{\nu}^{2},$ for the electron neutrino, which were due to systematic effects, and not a true $m_{\nu}^{2}<0.$ The KATRIN experiment is the latest and most accurate experiment to measure $m_{\nu}^{2}$ and it has corrected for many of those earlier systematic effects, especially molecular final states.
+(Schneidewind2024, ) Thus, one assumption in direct mass experiments is that our current understanding of the molecular final states is accurately known.
+Would a future statistically significant $m^{2}<0$ mass then be attributed to a residual unknown systematic effect rather than an “unphysical” $m_{\nu}^{2}<0$ value?
+The use of the term unphysical for $m_{\nu}^{2}<0$ neutrinos in KATRIN publications suggests the former is more likely.
+(Drexlin2026, )
+As of 2024 KATRIN has found only an upper limit to the mass, $m_{\nu}<0.45eV/c^{2}$ and by the end of the experiment they expect to reach a final sensitivity of better than $0.3eV/c^{2}$ at $90\%$ C.L.
+In addition, the best fit value as of 2024 is $m^{2}=-0.14^{+0.13}_{-0.15}eV^{2}.$ (KATRIN2024, ) This result, of course, leaves the true sign of $m^{2}$ ambiguous given the uncertainty.
+It is an interesting coincidence that this KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}.$ Ehrlich2015 While the statistical uncertainty on the $m^{2}$ value will be reduced as KATRIN takes more data, it might seem virtually certain that it will not shrink enough to eventually reveal its true sign to a $90\%$ C.L.
+It is routinely claimed by KATRIN collaborators that direct mass experiments are “model-independent” and they make no assumptions (aside from energy conservation).
+(Drexlin2013, ; KATRIN2022, ; Parno2025, ) However, in addition to the several assumptions already noted, there is one other assumption routinely made.
+Namely, it is assumed that if the electron neutrino effective mass is close to zero, then the three masses $m_{j}$ contributing to it in Eq.
+1 must also be close to zero, and hence not be experimentally resolvable.
+This assumption means that it is fine to fit the spectrum using a single mass.
+However, if one of the three masses satisfies $m_{j}^{2}<0,$ Eq.
+1 would allow the three effective masses for $\nu_{e},\nu_{\mu}$ and $\nu_{\tau}$ to all be very close to zero, even when the individual $m_{j}$ are not.
+In fact, it will be shown using published neutrino data for SN 1987A that the three $m_{j}$ comprising the electron neutrino apparently have surprisingly large values.
 I.4 Supernova SN 1987A neutrinos
 Three neutrino detectors operated at the time supernova SN1987A occurred at 7:35 UT on Feb 23, 1987.
-Those detectors (Kamiokande II, IMB, and Baksan) recorded bursts of 12, 8 and 5 antineutrinos respectively at approximately the same time, and they recorded the visible energy, incoming direction and arrival times of the neutrinos.
+Those detectors (Kamiokande II, IMB, and Baksan) recorded “bursts” or seconds-long groupings of 12, 8 and 5 antineutrinos respectively at approximately the same time, and they recorded the visible energy, incoming direction and arrival times of the neutrinos.
 The three detectors were not perfectly synchronized, with a time difference between them being perhaps a few seconds.
-This lack of synchronization is usually accounted for by assuming the first event in each detector occurred at time $t=0.$ The arriving electron antineutrinos were detected based on inverse beta decay
-$\bar{\nu_{e}}+p\rightarrow n+e^{+}$ (1)
-so that the antineutrino energy was actually 1.3 MeV higher than the recorded visible $(e^{+})$ energy.
+This lack of synchronization is usually accounted for by assuming the first event in each detector occurred at time $t=0.$ The arriving electron antineutrinos were detected based on inverse beta decay $\bar{\nu_{e}}+p\rightarrow n+e^{+}$ so that the antineutrino energy was actually 1.3 MeV higher than the recorded visible $(e^{+})$ energy.
 Those three neutrino bursts occurred about three hours before the star emitting them suddenly brightened.
@@
 I.5 The 5 hour early LSD neutrino burst
-In addition to the three detectors, there was a fourth small detector known as LSD ln 1987 which was located under Mont Blanc.
-This detector was in fact better shielded from background radiation than the other three, and had a lower energy threshold, so in these two respects was more sensitive.
+In addition to the three detectors, there was a fourth small detector known as LSD which was located under Mont Blanc.
+This detector was in fact better shielded from background radiation than the other three, and had a lower energy threshold, so in these two respects it was more sensitive.
 The LSD detector reported a burst of five neutrinos which occurred 484 min before the other three detectors’ received bursts.
@@
 However, that can also be explained because the small LSD detector did observe two neutrinos that time (not enough to qualify as a burst).
-Two final major problem with the LSD burst being associated with SN 1987A are (1) the very low energies of all the 5 neutrinos comprising it, and (2 ) their nearly equal energies or the virtual monochromaticity of the burst.
+Two final major problems with the LSD burst being associated with SN 1987A are (1) the very low energies of all the 5 neutrinos comprising it, and (2 ) their nearly equal energies or the virtual monochromaticity of the burst.
 In what follows we accept that the LSD neutrino burst was associated with SN 1987A, because all five objections to it are answerable, but we address the monochromaticity problem later.
@@
 It is however possible that the data points will all be found to be associated with only one or two straight lines rather than three lines if two or three of the lines (or neutrino masses) are too close to distinguish from one another.
-III Fitting 30 data points to 3 lines
-Thirty neutrinos is not a very large number, but the distinctiveness of a pattern counts for far more than the sheer number of events.
+III Fitting 37 data points to 3 lines
+Thirty-seven neutrinos is not a very large number, but the distinctiveness of a pattern counts for far more than the sheer number of events.
 Thus, the $\Omega^{-}$ particle which led to quarks was confirmed in 1964 at Brookhaven National Laboratory by observing just one single“Gold Standard” event.
-(Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the 30 neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only one data point very near the origin is within an error bar width of two lines.
-The least square fit of the 30 data points to three lines through the origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig.
-1 yielding 27 degrees of freedom.
+(Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only one data point very near the origin is within an error bar width of two lines.
+The least squares fit of 37 data points to three lines through the origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig.
+1 yielding 34 degrees of freedom.
 The parameter H, the size of the constant horizontal error bar on each data point in Fig.1 is also adjusted.
@@
 If a data point lies close to two straight lines, we simply choose the line that gives the smaller contribution to chi square.
-Only a few of the data points lie beyond $1\sigma$ of a straight line.
-The good fit shown in Fig 1 has a chi square of 25.4.
-Given 27 degrees of freedom we find $p=55\%.$ This excellent best fit was for a horizontal error bar of two seconds on all data points, and the fit would still be acceptable but significantly worse for a one second horizontal error bar ($p=15\%$.) For a fuller description of how the $\chi^{2}$ fit probability depends on the size of the horizontal error bar see Fig.
-2 We have seen that exactly simultaneous neutrino emissions requires the data must lie on one of three lines through the origin in a $1/E^{2}$ versus t plot.
-Therefore the goodness of fits using 1 and 2 sec horizontal error bars shows they will continue to lie on or near these lines if the time window of neutrino emissions is also no more than 1-2 seconds.
-Figure 1: Best fit of 30 data points $(t,1/E^{2}),$ one for each neutrino to one of three straight lines that pass through the origin.
+III.1 The best fit
+The best fit of 37 data points $(t,1/E^{2}),$ one for each neutrino to one of three straight lines that pass through the origin is shown in Fig.
+1.
 Error bars in both horizontal and vertical are included in the fit.
-The fit has $\chi^{2}=25.4$ for 27 dof, which yields $p=55\%$.
 The three best fit masses $m_{j}$ are obtained from the three best fit slopes according to Eq.
@@
 Note the different scales used for $t>0$ and $t<0$ The five LSD data points at t=17ksect=-17ksec have been slightly separated from each other for clarity.
+The hard to see data point at t=2.5t=2.5sec carries a great deal of weight because of the small vertical error bar.The good fit shown in Fig 1 has a chi square of 29.9.
+Given 34 degrees of freedom we find $p=67\%.$ This good fit was for a horizontal error bar of $\pm 0.5$ seconds on all data points, and the fit would be much better for a $\pm 1.0$ second horizontal error bar ($p=91\%$.) For a fuller description of how the $\chi^{2}$ fit probability depends on the size of the horizontal error bar see Fig.
+2 We have seen that exactly simultaneous neutrino emissions requires the data must lie on one of three lines through the origin in a $1/E^{2}$ versus t plot.
+Therefore the goodness of fits using 0.5 and 1 sec horizontal error bars shows they will continue to lie on or near these lines if the time window of neutrino emissions is also around 0.5 to 1 seconds.
+Figure 1: Best fit of 37 data points $(t,1/E^{2}),$ one for each neutrino to one of three straight lines that pass through the origin.
+See text and Table I for the meaning of the shaded region containing 7 “extra” Kamiokande II data points.
+Seven data points in shaded region in Fig.1
+time (sec) Hits     E(MeV)     $\Delta E$ (MeV)
+13.7 30 10.8 2.1
+15.1 31 11.2 2.1
+15.4 25 9.0 1.9
+16.5 26 9.4 1.9
+18.9 32 11.5 2.1
+18.9 28 10.1 2.0
+22.8 33 11.9 2.2
+Table 1: Seven “extra” Kamiokande II data points extracted from Fig.
+4f of Hirata et al.
+(Hirata1988, ).
+Thse seven data points have $N_{hits}>24$ and $13<t<25$ seconds.
+The energy for each data point was estimated based on: $E=0.36N_{hits}.$ The uncertainty in the energy (before 1988) was $\Delta E/E=22\%/\sqrt{E/10MeV}$ (Hirata1989, )
+III.2 The 7 extra Kamiokande II neutrinos
+One problem in fitting the SN 1987A data to three straight lines is deciding what is the maximum time to use in the fit.
+This is important since after many seconds neutrinos are increasingly likely to be just background and not be associated with SN 1987A.
+Ref.
+(Hirata1988, ) for Kamiokande II arbitrarily chose $t=12$ sec as the time cutoff.
+This choice resulting in their tabulating data for 12 SN 1987A neutrinos.
+It was concerning to the author, however, that the last data point at $t=12$ seconds in Fig.
+1 is $1.9\sigma$ above the fitted line.
+This concern motivated the desire to add more consecutive data points beyond $t=12$ sec, and seven of these were extracted by the author from ref.
+(Hirata1988, ) as explained in Table I.
+These data points are very unlikely to be due to background which lies below 20 hits, i.e.
+20 photomultipliers activated, and adding them has had little effect on the goodness of the fit to three straight lines or their slopes.
 Based on Eq.
 5 the three masses corresponding to the slopes of the three straight lines in Fig.
-1 are $m_{1}=22.4eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ for the $m^{2}<0$ line.
-The author first discovered this three line fit in 2013, when he interpreted the result in terms of a 3 + 3 model (three pairs of active-sterile neutrinos – see Fig.3.) The big difference from the author’s previously published values of this fit (Ehrlich2013, ; Ehrlich2019, ) is in the value of $m_{2}$ because of the past omission of horizontal error bars.
-Those error bars are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line.
-We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.70eV/c^{2}$ and the $2.70eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration (both equal to within $1\%$), as well as its non-observation by KATRIN, (Serebrov2024, ; KATRIN2025, ) a matter discussed at length in section IV.10.
-III.1 Fixing a limitation of sterile neutrino searches
-Of course, the existence of sterile neutrinos is highly controversial, there are probably more negative search results than positive ones.
+1 are $m_{1}=21.6eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-192,000eV^{2}/c^{4}$ for the $m^{2}<0$ line.
+The author first discovered this three line fit to the SN 1987A data in 2013, when he interpreted the result in terms of a 3 + 3 model (three pairs of active-sterile neutrinos – see Fig.3.) There are two significant differences from the author’s previously published work on this fit (Ehrlich2013, ; Ehrlich2019, ).
+The first difference is in the use of data for Kamiokande II beyond the 12 neutrinos they tabulated, which slightly changes the value of $m_{1}.$ The second major difference is in the change in the value of $m_{2}$ because of the past omission of horizontal error bars.
+Those horizontal error bars are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line.
+We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.70eV/c^{2}$ and the $2.70eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration, as well as its non-observation by KATRIN, (Serebrov2024, ; KATRIN2025, ) a matter discussed at length in section IV.10.
+III.3 An omission in most sterile neutrino searches
+The existence of sterile neutrinos is highly controversial, and there are probably more negative search results than positive ones.
 Thus, neither MicroBooNE, nor IceCube, nor KATRIN nor DayaBay report any evidence for them.
 (Micro2025, ; Aartsen2020, ; KATRIN2025, ; An2024, ).
-Nevertheless, an omission in these and other experiments is that they have not sought evidence for the presence of multiple sterile neutrinos simultaneously.
-In effect, if you are only searching for one sterile neutrino when two or three are present they can mask each others presence.
-One way to avoid this problem is to search for two or three in the same fit to the data, and another way is to find a way to spatially separate any sterile neutrinos based on their masses, or find a place where nature causes this separation for us – see next section.
-III.2 More evidence for the two $m^{2}>0$ masses
-It is believed that dark matter holds together both our galaxy as well as clusters of galaxies.
-One would expect that if both heavy and light $m^{2}>0$ sterile neutrinos were present throughout space the heavy ones would preferentially congregate in an individual galaxy and the lighter ones would preferentially congregate an a galaxy cluster.
-M.
+Nevertheless, an omission in these and other experiments is that most have not sought evidence for the presence of multiple sterile neutrinos simultaneously.
+There are several possible remedies for this omission, the simplest being to search for two or three sterile neutrinos in the same fit to the data, and another being to find a place where nature causes a spatial separation between light and heavy sterile neutrinos.
+To date the only search for the pair of $m^{2}>0$ sterile neutrinos suggested by the SN 1987A data has been carried out by M.
 H.
-Chan and the author have carried out fits to the “observed” dark matter distributions in both the Milky Way galaxy and in four clusters of galaxies.
-(Chan2014, ) The goal was to see what mass sterile neutrino would match the observed profile in each case.
-This was done by deriving an equation of state for the sterile neutrinos dependent on their mass, and then finding the dark matter radial profile in both the Milky Way and four groups of galaxies.
+Chan and the author.
+(Chan2014, ) Essentially, we first derived an equation of state for the sterile neutrinos dependent on their mass, and then separately found the dark matter radial profile in both the Milky Way and four groups of galaxies.
 Finally, the mass was varied to find the value that best fit the observed profile in each case.
-The results of these calculations showed that the two $m^{2}>0$ values from the 3-mass SN 1987A fit would do nicely with the heavier mass holding clusters together as would be expected.
-It seems unlikely one could find another case where nature is going to be so kind as to separate spatially different mass dark matter particles.
-However, as we shall see in section IV.10, KATRIN has a way to search for all three neutrino masses including the tachyonic one simultaneously in a single fit.
-Figure 2: The probability from the least squares fit to three straight lines through the origin versus the size of the horizontal error bars, assumed to be the same for all 30 data points.
-Note that there is no “best” fit to three straight lines unless one first fixes the size of the horizontal error bars because for $H>>5$ sec error bars the fit probability $p\approx 100\%$
-Figure 3: $3+3$ neutrino mass squared diagram postulated by the author in 2013, with updated $m_{1}$ and $m_{2}$ values.
-There are 3 active-sterile doublets (solid and dotted lines) split by three $\Delta m^{2}$ The three $\Delta m^{2}$ are for solar, atmospheric and short baseline (sbl) oscillations.
+The results of these fits showed that the two $m^{2}>0$ values from the 3-mass SN 1987A fit would indeed fit the galaxy clusters (lighter mass) and the Milky Way (heavier mass).
+This result offers a confirmation of the $m_{1}=21.6eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ masses deduced from the SN 1987A three mass fit.
+Figure 2: The probability from the least squares fit to three straight lines through the origin versus the size of the horizontal error bars, assumed to be the same for all 37 data points.
+Note that there is no “best” fit to three straight lines unless one first fixes the size of the horizontal error bars because for $H>1.5$ sec error bars the fit probability $p\approx 100\%$
+Figure 3: $3+3$ neutrino mass squared diagram postulated by the author in 2013, with updated $m^{2}$ values.
+There are 3 active-sterile doublets split by three $\Delta m^{2}$ for solar, atmospheric and short baseline (sbl) oscillations.
 Diagram is not to scale.
@@
 Many physicists regard $m^{2}<0$ tachyons as ghost particles.
-Illustration of a ghostly nightmare is from Shutterstock.
+Illustration from Shutterstock.
 IV Discussion
-In this section we discuss 10 issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0,$ really does imply we can deduce the three neutrino masses from the SN 1987A data, and it is not just a mirage or background fluctuation.
+In this section we discuss 12 issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0$ mass, really does imply we can deduce the three neutrino masses from the SN 1987A data, and it is not just a mirage or background fluctuation.
 IV.1 Why not exclude the LSD data?
@@
 This rejection was not because it provided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous.
-However since this issue and all the other reasons many people mistakenly have for rejecting the LSD data can be answered, there is no valid reason not to keep LSD data, especially because with three active neutrinos the mandated number of straight lines about which clustering should occur is three.
+However since this issue and all the other reasons many people have for rejecting the LSD data can be answered, there is no valid reason not to keep LSD data, especially because with three active neutrinos the mandated number of straight lines about which clustering should occur is three.
+The burden of proof here should not be why not exclude this burst, but rather why not keep it – and there really is no compelling reason to ignore it.
 IV.2 Why are the LSD data so bunched in time?
@@
 The three straight line result would be more believable perhaps if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line.
-But recall that the $t<0$ data has a kilosecond time scale.
+But recall that the $t<0$ data have a kilosecond time scale.
 If the five LSD data points were spread out along the negative sloped line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst.
@@
 Many phenomena in nature on the fringes of the known have alternative explanations.
-All the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems.
+Most of the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems.
 Thus, perhaps the most likely alternate possibility is that a supernova can bang more than once.
 In fact, Galeotti and Pizzella have even claimed there was a third bang– not in the LSD detector but in Kamiokande II at a time about 20 min after the main burst.
-(Galeotti2007, ) However, Kamiokande II does not acknowledge this burst.
-Supernovae undergoing multiple bangs is a known phenomenon (known as Pulsational pair-instability supernovae), which can occur for very massive stars according to Woosley.
-(Woosley2017, ) According to him SN 1987A had a helium core of 5 to 6 $M_{sun}$ which is characteristic of an $18-20M_{sun}$ main sequence star.
-However, this mass is too small to experience the pair instability which requires He core masses of at least $37M_{sun}.$ Additionally, Woosley notes that pulsational pair instaility supernovae do not produce detectable neutrino bursts.
-Among the other alternative explanations of the LSD burst are:
+(Galeotti2007, ) Supernovae undergoing multiple bangs is a known phenomenon (known as Pulsational pair-instability supernovae), which can occur for very massive stars according to Woosley.
+(Woosley2017, ) However, the mass of SN 1987A was too small to experience the pair instability and in any case the phenomenon does not produce detectable neutrino bursts.
+Other explanations of the LSD burst include:
   • If it was a statistical fluctuation it was a really big one $p=1.4\times 10^{-6}.$ In fact among the LSD signal candidates for the detection of a neutrino burst selected over 14 years of operation, it has the lowest background imitation probability.
-  • If it was a genuine, low-energy pulse (e.g., from neutron capture on iron slabs around the detector), simulations show that their number is completely negligible for all of the neutrino radiation parameters considered.
-(Manukovskiy2022, ).
-  • If it was due to 8 MeV dark matter as suggested in ref.
-(Ehrlich2021, ) this would still create the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD neutrinos falling on the $m^{2}<0$ line.
-  • The only other explanation known to this author for the LSD burst besides tachyons is the Froggatt-Nielsen proposal of “dark matter balls” inside the the stellar core (Froggatt2015, ).
+  • If it was a genuine, low-energy pulse of neutrinos (e.g., from neutron capture on iron slabs around the detector), simulations show that their number is negligible. (Manukovskiy2022, ).
+  • If it was due to 8 MeV dark matter as suggested in ref. (Ehrlich2021, ) this would still create the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD burst.
+  • The only other explanation known to this author for the LSD burst besides tachyons is the Froggatt-Nielsen proposal of “dark matter balls” inside the stellar core (Froggatt2015, ).
 IV.5 Best size of the horizontal error bars?
@@
 Clearly from Fig.
-2 it is about 1-2 seconds.
+2 it is about 0.5 to 1.0 seconds.
 IV.6 The time window for neutrino emissions?
@@
 Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background.
-These results all buttress our case that the SN 1987A neutrinos were mostly emitted during a 1-2 sec time window, and that our use of 2 sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified.
-A 1 to 2 sec time window is apparently small enough not to badly degrade the fit of 30 neutrino data points to three straight lines.
+These results all buttress our case that the SN 1987A neutrinos were mostly emitted during a 0.5 to 1.0 sec time window, and that our use of $\pm 0.5$ sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified.
+A 0.5 to 1.0 sec time window is apparently small enough not to badly degrade the fit of 37 neutrino data points to three straight lines.
 Figure 4: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al.
@@
 Clearly, they still would lie on three lines if emitted simultaneously, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics.
-However, we would not be able to find these lines because for 3 small masses they would lie on very many almost vertical lines filling the time window – see the 2 second error bar in Fig, 5.
+However, we would not be able to find these lines because for 3 small masses they would lie on very many almost vertical lines filling the time window in Fig.
+5.
 In effect each point lies on its own vertical line filling the time window.
-This is obviously far from what is the case for the actual data, which suggests that the neutrino mass does not satusfy $m<<1eV/c^{2}.$
-Figure 5: How 50 neutrinos might appear on a $1/E^{2}$ versus t plot if their mass $m<<1eV/c^{2}$ and they were emitted during a time window of 2 seconds.
+This is obviously far from what is the case for the actual data, which suggests that the neutrino mass does not satisfy $m<<1eV/c^{2}.$
+Figure 5: How 50 neutrinos might appear on a $1/E^{2}$ versus t plot if their mass $m<<1eV/c^{2}$ and they were emitted during a time window of under 1 second.
 No congregation of the points about specific lines would be observed, completely unlike the real SN 1987A data which clearly clusters along three straight lines.
-Essentially, each of the 50 points lies on its own different vertical line somewhere within the 2 second time window and no two points lie on the same line.
-IV.8 Consistency with cosmological data?
+Essentially, each of the 50 points lies on its own vertical line somewhere within the 1 second time window and no two points lie on the same line.
+IV.8 The $\sum m_{\nu}$ constraint: a mass or flavor sum?
 An apparent conflict exists between the three masses from the SN 1987A data fit and cosmological data.
-That conflict on the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ from cosmological data can be avoided, if one of the tachyon masses is negative gravitationally.
+That conflict with the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ can be avoided, if one of the tachyon masses is negative gravitationally.
 (Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, meaning they have a repulsive gravitational effect.
 The current upper limit on the total sum of neutrino masses from cosmology has tightened to $\sum m_{\nu}<0.07-0.12eV/c^{2}$ (Shao2025, ), but some analyses of the data even favor a negative value for the sum of the neutrino masses.
-(Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found in the SN 1987A data, which could be avoided if the mass sum were over flavor state masses not mass states.
+(Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found in the SN 1987A data, which could be avoided provided the mass sum were over flavor state masses not mass states.
 In the recent literature the issue of whether the sum is over mass or flavor states is not widely discussed.
@@
 Cosmological neutrinos were probably born in the universe’s infancy, decoupling from matter and radiation when the universe was only about 1 second old.
-At that time the two dominant processes capable of producing neutrinos was $Z^{0}$ decay or $e^{+}e^{-}$ collisions into neutrino antineutrino pairs, and would both produce equal numbers for each flavor.
+At that time the two dominant processes capable of producing neutrinos were $Z^{0}$ decay or $e^{+}e^{-}$ collisions into neutrino antineutrino pairs, and both reactions would produce equal numbers of neutrinos for each flavor.
 For example, $Z^{0}$ decays would result in equal $6.7\%$ branching ratios for each flavor, thus creating an equal number density N of each.
-Likewise for $e^{+}e^{-}$ collisions.
 Thus, the cosmological neutrino mass/energy density produced in that manner would be:
@@
 1 seem preposterously large, and they would appear to be inconsistent with data from the KATRIN experiment implying $m_{\nu}<0.45eV/c^{2}$.
-As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the electron neutrino effective mass has $m_{k}^{2}<0$ the effective mass, as normally defined, $m_{e}^{2}=\sum|U_{e,k}|^{2}m_{k}^{2}$ can be arbitrarily close to zero or even negative, even when the $m_{k}^{2}$ are all very large, as in the SN 1987 fit.
-KATRIN has further crucial relevance quite apart from the limit it imposes by its measured $\nu_{e}$ effective mass.
-IV.10 A missing $m_{4}=2.70eV/c^{2}$ sterile neutrino?
+As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the electron neutrino effective mass has $m_{k}^{2}<0$ the effective mass (see Eq.1) can be arbitrarily close to zero or even negative, even when the $m_{k}^{2}$ are all very large, as in the SN 1987 best fit.
+In fact, given that the $3\times 3$ unitary mixing matrix has nine free parameters, all three effective masses could be close to zero.
+IV.10 A “missing” $m_{4}=2.70eV/c^{2}$ sterile neutrino?
 KATRIN recently reported finding no evidence for sterile neutrinos in its data based on the observed shape of the tritium beta decay spectrum.
 (KATRIN2025, ).
-It conducted this search by fitting its data to a sum of two spectra: that for a zero mass neutrino plus the spectrum for a sterile neutrino of mass $m_{4}$, which was assumed to take on various masses and various weights.
-KATRIN then used a logarithmic grid of trial $m^{2}$ and $sin^{2}2\theta$ values that spans from $m^{2}=0.1eV^{2}$ to $1600eV^{2}$ and weight $sin^{2}2\theta$ from 0.001 to 0.5.
-At each grid point, the $\chi^{2}$ function is minimized with respect to three free parameters.
-One can then see how the goodness of fit varies over the grid, and how much of the grid is excluded based on a bad fit.
+It conducted this search by fitting its data to a sum of two spectra: that for a zero mass neutrino plus the spectrum for a sterile neutrino of mass $m_{4}$, which was assumed to take on various values and various weights $\alpha=sin^{2}2\theta.$ At each $(m_{4},\alpha)$ grid point, the $\chi^{2}$ function is minimized with respect to three free parameters.
 By examining the excluded regions of the plane based on the high $\chi^{2}$ KATRIN concluded there was no evidence for sterile neutrinos at a high confidence level in their data over the entire grid.
-This negative result was surprising because the Neutrino-4 group had previously found evidence for a $2.70eV/c^{2}$ sterile neutrino in their neutrino oscillation experiment.
-Moreover, the KATRIN group is always very careful in their analysis of their data, so missing a signature of a sterile neutrino seems unlikely.
+This negative result was was in conflict with the Neutrino-4 group that had previously found evidence for a $2.70eV/c^{2}$ sterile neutrino in their oscillation experiment.
 Moreover, the conflict of KATRIN with the Neutrino-4 experiment was not a small one, since the data of the GALLEX, SAGE, and BEST experiments confirms the parameters of neutrino oscillations $\Delta m^{2}_{14}=7.3eV^{2}$ (or $m_{4}=2.70eV/c^{2}$ and $\sin^{2}2\theta_{14}=0.36)$ and the confidence level of the Neutrino-4 sterile neutrinos combined with the data of the GALLEX, SAGE, and BEST experiments is now $5.8\sigma.$ (Serebrov2024, )
 IV.11 “The dog that didn’t bark” (the key section!)
-It is possible that KATRIN failed to see any sterile neutrinos because they simply do not exist, despite the above high confidence kevel, but here we consider a much more interesting possibility.
+It is possible that KATRIN failed to see any sterile neutrinos because they simply do not exist, but here we consider a much more interesting possibility.
 Neutrino - 4 and the other three experiments that saw the $2.70eV/c^{2}$ neutrino were doing a short baseline disappearance oscillation experiment, which observed the disappearance probability $P(\bar{\nu_{e}}\rightarrow\bar{\nu_{e}}).$ This type of experiment would work just as well if the disappearing neutrinos were disappearing into sterile or active neutrinos.
-However, KATRIN’s search was specifically looking for sterile neutrinos and it would not work for active ones as conducted.
+However, KATRIN’s search was specifically looking for sterile neutrinos and it would not have worked for active ones as it was conducted.
 (KATRIN was not being careless here, they just assumed like everyone else that neutrinos with large mass would have to be sterile) Of course, if the main claim of this paper is correct, and if the $2.70eV/c^{2}$ neutrino is an active neutrino (as well as a sterile one - see Fig.
-4), then KATRIN’s failure to see it is understandable because their search method just would not work for active neutrinos.
+3), then KATRIN’s failure to see it is understandable because their search method was not appropriate for active neutrinos.
 Specifically, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution.
-(Why not you ask?
-Because the zero mass effective mass is made up of the three specific masses we found including one with $m^{2}<0$) Rather, they needed to fit their data using a combined differential spectrum $R_{C}$ for.the three masses $m_{1}=2.70eV/c^{2},m_{2}=22.4eV/c^{2}$ and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ with three unknown adjustible weights, $\alpha_{j},$ that is:
-$R_{C}=\sum\alpha_{j}R_{\beta}(E,E_{0},m^{2}_{\nu,j})$ (7)
-where $R_{\beta}$ is the usual beta decay formula for a given $m^{2}$, E is the electron energy and $E_{0}$ is the fitted endpoint energy for tritium beta decay.
+That’s because the zero mass effective mass is made up of the three specific masses we found for SN 1987A including one with $m^{2}<0.$ So, they should have fit their data using a combined differential spectrum $R_{C}$ for the three masses $m_{1}=2.70eV/c^{2},m_{2}=21.6eV/c^{2}$ and $m_{3}^{2}=-192,000eV^{2}/c^{4}$ with three adjustible weights, $\alpha_{j},$ that is:
+$R_{C}=\sum\alpha_{j}R_{\beta}(E,E_{0},m^{2}_{\nu,j})+R_{Bkgd}$ (7)
+where $R_{Bkgd}$ is the constant background rate, $R_{\beta}$ is the usual beta decay formula for a given $m^{2}$, E is the electron energy and $E_{0}$ is the fitted endpoint energy for tritium beta decay.
 For the weights $\alpha_{j}$ only two are independent since $\sum\alpha_{j}=1,$ and the $\alpha_{3}$ for the tachyonic mass is obviously very small in view of the very large magnitude of $|m_{3}^{2}|.$
 To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the observed spectrum are found.
-With the three masses set to their specified SN 1987A values, KATRIN should ideally observe a small concentrated area where good fits are found surrounded by a large area of poor fits.
+With the three masses set to their specified SN 1987A values (or else used as three pull terms in the fit), KATRIN should ideally observe a small concentrated area on a $(\alpha_{1},\alpha_{2})$ grid where good fits are found surrounded by a large area of poor fits.
 They could also do the same using other masses, where no good fits should be found.
-The best fit on the $\alpha_{1}$ and $\alpha_{2}$ plane should yield the effective mass from $m^{2}=\sum\alpha_{j}m^{2}_{j},$ and it should be a very small negative value.
+The best fit on the $\alpha_{1}$ and $\alpha_{2}$ plane should also yield the effective mass from $m^{2}=\sum\alpha_{j}m^{2}_{j},$ and it should be a very small negative value.
 Finally, for the functional form of the spectrum for the tachyonic third mass, $m_{3}^{2},$ KATRIN could just use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it.
@@
 In summary, if KATRIN were to redo its analysis with the preceding suggested changes it could be a decisive test of whether the SN 1989A data really can provide the three neutrino mass values, including the tachyonic one $m_{3}$.
-A positive KATRIN result which confirms all three masses from the SN 1987A data fit would be of some considerable interest.
 The missing $2.70eV/c^{2}$ sterile neutrinos in the KATRIN search was the key clue here.
@@
 IV.12 Multiple sterile neutrinos in oscillation data?
-Anatolii Serebrov and his Neutrino 4 team, have argued on cosmological grounds for there being two heavier sterile neutrinos in addition to the one found at $m_{1}^{2}=2.70^{2}eV^{2}/c^{4},$ which of course was also earlier inferred from the SN 1987A data analysis.
-Assuming a second sterile neutrino of mass $m_{2}^{2}=22.4eV^{2}/c^{2},$ as mandated by that data, the oscillation wavelength would be 69 times shorter than the $m_{1}$ sterile neutrinos.
-It is highly unlikely that such short wavelength oscillations are detectable in the data of Neutrino 4, but if they were one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find the likelihood probability of the fit to the data for that pair of mass values.
-One could then see where on the plane that likelihood maximizes.
-If there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},22.4^{2})eV^{2}/c^{2},$ that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data.
-You might think that the most promising experiments to look at would be the four data sets identified by Serebrov in ref.
-(Serebrov2024, ), which already have shown evidence of the $2.70eV/c^{2}$ mass at a cumulative $5.8\sigma$.
-Unfortunately, however, the very short wavelength of $22.4^{2}eV^{2}/c^{2}$ neutrinos almost certainly make these oscillations undetectable in existing short baseline reactor data.
-There may be over 20 projects in the works exploring the ultra-short baseline region ($L<10$ meters) in reactor experiments, (Serebrov2021, ) but no experiments known to the author could detect stable sterile neutrinos of such a high mass, with the exception of IceCube.
-In fact, IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in seven years of data.
-(Abbasi2024, ) Even though they reported a negative result their best fit did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass whuch was strongly preferred over the no oscillation solution by 10.7 to one.
-Furthermore, the IceCube energy range (up to $E=100TeV$) would make them sensitive to sterile mass values as high as $m_{2}^{2}=22.4^{2}eV^{2}/c^{4}.$ Thus, it might make sense for them to rexamine their data for evidence of a double oscillation for those two specific masses.
-Making such a search even more promissing, IceCube has already begun to explore the region of $\Delta m^{2}>10eV^{2}/c^{4},$ and found a preference for $3+1$ oscillations.
-(Biennow2018, ) A search using the method outlined for the Neutrino-4 data might prove particularly promising.
-Although the SN 1987A data analysis specified three neutrino masses it seems highly unlikely that the $m_{3}^{2}=-200,000eV^{2}/c^{4}$ neutrino would leave any detectable signal in IceCube or any other oscillation data (unlike KATRIN where it should be detectable).
+The need for doing searches for multiple sterile neutrinos in a single search has already been noted in section III.3.
+How would one go about seeking evidence for both $m_{1}^{2}=2.70^{2}eV^{2}/c^{4},$ and $m_{2}=21.6eV^{2}/c^{4}$ in oscillation data, bearing in mind that the oscillation wavelength for $m_{2}$ would be $21.6^{2}/2.7^{2}=64$ times shorter than the $m_{1}$ sterile neutrinos?
+It is unlikely that such short wavelength oscillations are detectable in the data of Neutrino 4, but if they were, one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find the likelihood probability of the fit to the data for that pair of mass values.
+One could then see if there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},21.6^{2})eV^{2}/c^{2}.$ A more likely possibility would be the IceCube experiment.
+IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in 10.7 years of data.
+(Abbasi2025, ) Even though IceCube reported a negative result their best fit they did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass which was strongly preferred over the no oscillation solution by 40 to one $p=2.5\%.$
 V Summary and the ultimate test
-This paper has made six claims that many readers may understandably consider very difficult to believe – see Table 1.
+This paper has made five claims that many readers may understandably consider very difficult to believe – see Table 2.
 Some of the claims will likely not be accepted as true without the strongest corroborating evidence.
-After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data, that might be enough.
-This might also be the case if oscillation experiments were to find such evidence doing the analysis mentioned in the previous section.
+After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data, that might convince many skeptics.
 1. The LSD burst was due to SN 1987A but no 2nd bang
-2. Upper limits on the 3 neutrino $m^{2}_{\nu,j}$ can be avoided
-3. 3 $m_{j}^{2}$’s really can be found from 30 SN 1987A neutrinos
+2. 3 $m_{j}^{2}$’s really can be found from 37 SN 1987A neutrinos
 3. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
-5. The three $m^{2}=2.7^{2}eV^{2},22.4^{2}eV^{2}$ and $2\times 10^{5}eV^{2}$
-6. KATRIN, IceCube or other data likely can confirm #5
-Table 1: six claims made in this paper.
+4. The three $m^{2}=2.7^{2}eV^{2},21.6eV^{2}$ and $1.92\times 10^{5}eV^{2}$
+5. KATRIN and other data likely can confirm the above
+Table 2: Five claims made in this paper.
 However, the most convincing test would be a new galactic supernova.
@@
 This 88 year-old does not have a great chance of seeing it, but let’s be an optimist.
-When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other five claims.
+When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other four claims.
 Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors.
-These neutrinos will definitely either confirm or reject the six claims made.
+These neutrinos will definitely either confirm or reject the five claims made.
 Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy.

### v7

@@
 $m^{2}c^{4}=\frac{2}{TM}$ (5)
-Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line through the origin having negative slope.
+Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line through the origin lying in the $t<0$ region.
 To the extent that neutrinos all started out from SN 1987A at nearly the same instant, then since there are known to be three (and only three) non-sterile neutrino masses, every neutrino should lie on or close to one of three straight lines passing through the origin.
@@
 (Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only one data point very near the origin is within an error bar width of two lines.
-The least squares fit of 37 data points to three lines through the origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig.
+The least squares fit of 37 data points to three lines having a common $t=0$ origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig.
 1 yielding 34 degrees of freedom.
@@
 (Hirata1988, ).
-Thse seven data points have $N_{hits}>24$ and $13<t<25$ seconds.
-The energy for each data point was estimated based on: $E=0.36N_{hits}.$ The uncertainty in the energy (before 1988) was $\Delta E/E=22\%/\sqrt{E/10MeV}$ (Hirata1989, )
-III.2 The 7 extra Kamiokande II neutrinos
+These seven data points have $N_{hits}>24$ and $13<t<25$ seconds.
+The energy for each data point was obtained by the author based on the linear relation: $E=0.363N_{hits}$ to an accuracy of $5\%$ (Ehrlich2018, ).
+The uncertainty in the energy (before 1988) was $\Delta E/E=22\%/\sqrt{E/10MeV}$ (Hirata1989, )
+III.2 The 7 “extra” Kamiokande II neutrinos
 One problem in fitting the SN 1987A data to three straight lines is deciding what is the maximum time to use in the fit.
@@
 (Hirata1988, ) for Kamiokande II arbitrarily chose $t=12$ sec as the time cutoff.
-This choice resulting in their tabulating data for 12 SN 1987A neutrinos.
+This choice resulting in their tabulating data only for 12 SN 1987A neutrinos.
 It was concerning to the author, however, that the last data point at $t=12$ seconds in Fig.
@@
 These data points are very unlikely to be due to background which lies below 20 hits, i.e.
-20 photomultipliers activated, and adding them has had little effect on the goodness of the fit to three straight lines or their slopes.
+20 photomultipliers activated, and adding them improved the goodness of the fit to three straight lines.
 Based on Eq.
@@
 (Micro2025, ; Aartsen2020, ; KATRIN2025, ; An2024, ).
-Nevertheless, an omission in these and other experiments is that most have not sought evidence for the presence of multiple sterile neutrinos simultaneously.
-There are several possible remedies for this omission, the simplest being to search for two or three sterile neutrinos in the same fit to the data, and another being to find a place where nature causes a spatial separation between light and heavy sterile neutrinos.
+Nevertheless, an omission in these and other experiments is that most have not sought evidence for the presence of multiple sterile neutrinos simultaneously, i.e., they do $3+1$ and not $3+2$ fits to data.
+There are two possible remedies for this omission, the simplest being to search for two or three sterile neutrinos in the same fit, and another one being to find a place where nature causes a spatial separation between light and heavy sterile neutrinos.
 To date the only search for the pair of $m^{2}>0$ sterile neutrinos suggested by the SN 1987A data has been carried out by M.
@@
 Chan and the author.
-(Chan2014, ) Essentially, we first derived an equation of state for the sterile neutrinos dependent on their mass, and then separately found the dark matter radial profile in both the Milky Way and four groups of galaxies.
+(Chan2014, ) Sterile neutrinos are a natural possibility for the identity of dark matter in the galaxy and clusters of galaxies.
+Essentially, the authors first derived an equation of state for the sterile neutrinos dependent on their mass, and then separately found the dark matter radial profile in both the Milky Way and four groups of galaxies.
 Finally, the mass was varied to find the value that best fit the observed profile in each case.
@@
 But recall that the $t<0$ data have a kilosecond time scale.
-If the five LSD data points were spread out along the negative sloped line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst.
+If the five LSD data points were spread out along the LSD line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst.
+The fact that the 5 LSD neutrinos were recorded at -17 ksec does, of course, not mean they went backward in time, only that they travelled faster than light, assuming they started at $t=0$ along with all the other three neutrino bursts.
+Because of that bunching at the left end of the LSD line the uncertainty in the slope (and the value of $m^{2}$) is $18\%$ or $m^{2}=-192,000\pm 34,000eV^{2}/c^{4}$.
 IV.3 Why do the LSD neutrinos have the same E?
@@
 2 it is about 0.5 to 1.0 seconds.
-IV.6 The time window for neutrino emissions?
-We can estimate the actual size of the time window for neutrino emissions using a simulation.
+Note that the time any individual neutrino is recorded is, of course, known to much greater precision than 0.5 seconds.
+However, the meaning of H takes into account the uncertainty of when the burst of neutrinos this neutrino was part of occurred, which is primarily determined by the choice of $t=0$ for the first arriving beutrino in a detector.
+IV.6 A simulation of SN 1987A neutrino emissions
 In 2025 Bozza et al.
@@
 Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background.
-These results all buttress our case that the SN 1987A neutrinos were mostly emitted during a 0.5 to 1.0 sec time window, and that our use of $\pm 0.5$ sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified.
+These Bozza et al.
+results buttress our case that the SN 1987A neutrinos were mostly emitted during a 0.5 to 1.0 sec time window, and that our use of $\pm 0.5$ sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified.
 A 0.5 to 1.0 sec time window is apparently small enough not to badly degrade the fit of 37 neutrino data points to three straight lines.
+However, the use of as many as 9 free parameters in the Bozza spectral fit is a serious drawback, since with enough free parameters one can fit virtually any shape.
+In contrast, our fit to the $m^{2}>0$ neutrino spectrum used only two parameters, i.e., the two $m^{2}>0$ masses.
+Moreover, as long as the spread in neutrino emission times is small compared to the horizontal error bars used (as implied by Fig.
+2), relativistic kinematics requires that neutrinos of a given mass lie on straight lines having a specific slope, and there is no similar reqirement in the Bozza analysis.
 Figure 4: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al.
@@
 IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in 10.7 years of data.
-(Abbasi2025, ) Even though IceCube reported a negative result their best fit they did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass which was strongly preferred over the no oscillation solution by 40 to one $p=2.5\%.$
+(Abbasi2025, ) Even though IceCube reported a negative result their best fit they did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass which was strongly preferred over the no oscillation solution by 40 to one $p=2.5\%.$ To date, however, IceCube has only done a limited amount of searching for multiple sterile neutrinos.
+(Cabrera2024, )
 V Summary and the ultimate test

### v8

@@
-Can the 3 neutrino masses really be found using SN 1987A data?
+Finding the 3 neutrino masses using a one-phase $\nu$ emission model for SN 1987A
 Abstract
@@
 As a result, the author rediscovered what Ramanath Cowsik had found in 1988 about the neutrino masses being possibly revealed in the SN 1987A data.
-(Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) an invaluable clue from the KATRIN experiment’s reported failure to observe sterile neutrinos.
+(Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, (3) an extensive critique of the standard two-phase neutrino emission model of SN 1987A and (4) an invaluable clue from the KATRIN experiment’s reported failure to observe sterile neutrinos.
 (KATRIN2025, ).
-This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation or rejection of the preposterous-sounding claim suggested in the title.
-The paper also suggests a way that oscillation experiments and especially the IceCube experiment might supply such a confirmation.
+This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation of the preposterous-sounding claim suggested in the title.
 I.1 Are $m^{2}<0,$ and $v>c$ Tachyons “unphysical”?
@@
 In this section we discuss 12 issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0$ mass, really does imply we can deduce the three neutrino masses from the SN 1987A data, and it is not just a mirage or background fluctuation.
-IV.1 Why not exclude the LSD data?
-Initially, the author mistakenly rejected the LSD data.
+IV.1 Neutrino emission models: 1, 2 or 3 phases?
+Standard neutrino emission models for supernovae usually stipulate two phases: an intense brief accretion phase lasting less than a second followed by a $10+$ seconds long cooling phase.
+Often the accretion phase is divided into two, with a first part being an initial breakout burst lasting under 50 ms, as the shock wave breaks through the surface of the protoneutron star, resulting in a massive flash of electron neutrinos.
+For example, Bozza et al.
+have calculated the neutrino luminosity from SN 1987A using a two phase model.
+(Bozza2025, ) The functional form of their two-phase computed luminosity has 9 free parameters that are determined by a least squares fit to the SN 1987A data, and their computed luminosity is shown in Fig.
+4.
+Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec,$ and the bulk of the neutrino emissions occur for $t<1$ seconds (the accretion phase).
+In contrast to the Bozza et al.
+two-phase model, our fit to the SN 1987A data just involves a single brief neutrino accretion phase lasting at most 0.5 to 1.0 seconds, and no cooling phase.
+An emission lasting at most 0.5 to 1.0 seconds was seen to be small enough so as to qualify as ”near-simultaneeous” neutrino emissions, for the purpose of giving a good fit to three straight lines on a plot of $1/E^{2}$ versus time – one line for each neutrino mass.
+If two supernova neutrino emission models both give good fits to the SN 1987A data (the only test yet available), we may ask which model is more likely to be closest to reality, simplicity being an important criterion.
+A two phase emission model is less simple than a one phase model, and a three phase model would be even less simple.
+Actually, contrary to its description, the Bozza model in fact is a three phase model not two, as they admit in their paper: “the presence of a third stage is required in order to give a smooth behavior of the observable neutrino distributions in the whole time range.” (Bozza2025, ) Further favoring a one-phase model for SN 1987A over two or three is the fact that its cooling was peculiar and unexpected because it originated from a blue supergiant, rather than a red supergiant.
+This fact would lead to faster adiabatic cooling and a fainter initial object, which challenged traditional models.
+Giving still further support for a one phase model, the use of as many as 9 free parameters in the Bozza two-phase model is a serious drawback, since with that many free parameters one can find a good fit to virtually any shape emissions spectrum.
+In contrast, to the Bozza model, our one phase fit to the $t>0$ neutrino spectrum used only two parameters, namely the two $m^{2}>0$ neutrino masses found from the slopes in Fig.
+1.
+Thus, as long as the spread in neutrino emission times is small, relativistic kinematics requires that neutrinos of a given mass lie on straight lines having a specific slope.
+Of course, there is no similar basic physics requirement justifying the two phases of accretion and cooling.
+The only real justification for having those two separate phases is because that is what SN 1987A data seems to show, but only if the neutrino masses are small enough to ignore the differences in their energy-dependent travel times.
+However, this conclusion requires use of a circular argument, since their neutrino emission time spectrum only mirrors the observed arrival time spectrum if the neutrino masses are very small, and the Bozza model (unlike our one-phase model) just assumes this is the case.
+A final reason to be skeptical about the Bozza model is its choice of neutrino data to include in its analysis, particularly for the cooling region, i.e., $13<t<30$ seconds.
+Kamiokande II showed approximately 40 data points in this time interval having $N_{hits}>20$ $(E_{\nu}>7.2MeV)$ (Hirata1988, ).
+Of those 40 points spread over the 17 second interval Bozza selected only four data points for inclusion in their analysis, based on their neutrino scattering directions.
+(Bozza2025, ) Clearly, the goodness of the Bozza two (three) phase fit would depend greatly on their selection criteria on which of the 40 data points to include and exclude.
+Despite the advantages discussed in this section of the author’s single phase neutrino emissions model over the Bozza two (or three) phase model for $t>0$ neutrinos, his inclusion of the $t<0$ LSD neutrino burst undoubtedly will result in some skepticism.
+IV.2 Why not exclude the LSD data?
+Initially, the author did reject the LSD data.
 This rejection was not because it provided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous.
-However since this issue and all the other reasons many people have for rejecting the LSD data can be answered, there is no valid reason not to keep LSD data, especially because with three active neutrinos the mandated number of straight lines about which clustering should occur is three.
-The burden of proof here should not be why not exclude this burst, but rather why not keep it – and there really is no compelling reason to ignore it.
-IV.2 Why are the LSD data so bunched in time?
+However since this issue and most of the other reasons many people have for rejecting the LSD data can be dealt with, there is no valid reason not to keep LSD data, especially because given three active neutrinos, the mandated number of straight lines about which clustering of data points in Fig.
+1 should occur is three.
+The burden of proof here should not be why not exclude this LSD burst, but rather why not keep it – and there really is no compelling reason to ignore it.
+IV.3 Why are the LSD data so bunched in time?
 If the LSD burst of 5 neutrinos is genuine, it may look strange in Fig.
@@
 If the five LSD data points were spread out along the LSD line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst.
-The fact that the 5 LSD neutrinos were recorded at -17 ksec does, of course, not mean they went backward in time, only that they travelled faster than light, assuming they started at $t=0$ along with all the other three neutrino bursts.
-Because of that bunching at the left end of the LSD line the uncertainty in the slope (and the value of $m^{2}$) is $18\%$ or $m^{2}=-192,000\pm 34,000eV^{2}/c^{4}$.
-IV.3 Why do the LSD neutrinos have the same E?
+The fact that the 5 LSD neutrinos were recorded at -17 ksec does, of course, not mean they “went backward in time,” only that they travelled faster than light, assuming they started at $t=0$ along with all the other three neutrino bursts.
+IV.4 Why do the LSD neutrinos have the same E?
 The apparent near-constant energy of the LSD burst remains mysterious even to the physicists who accept its existence, but it is natural if the 5 neutrinos (and all others) must lie on one of three straight lines.
@@
 (Ehrlich2021, )
-IV.4 Alternate explanations for the LSD neutrinos?
+IV.5 Alternate explanations for the LSD neutrinos?
 Many phenomena in nature on the fringes of the known have alternative explanations.
@@
   • The only other explanation known to this author for the LSD burst besides tachyons is the Froggatt-Nielsen proposal of “dark matter balls” inside the stellar core (Froggatt2015, ).
-IV.5 Best size of the horizontal error bars?
+IV.6 Best size of the horizontal error bars?
 The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large for the method to work.
@@
 However, the meaning of H takes into account the uncertainty of when the burst of neutrinos this neutrino was part of occurred, which is primarily determined by the choice of $t=0$ for the first arriving beutrino in a detector.
-IV.6 A simulation of SN 1987A neutrino emissions
-In 2025 Bozza et al.
-have calculated the neutrino luminosity from SN 1987A using a new model they recently developed.
-(Bozza2025, ) They claim the analysis is in many ways more refined and accurate than previous ones with respect to the temporal structure of the neutrino luminosity.
-The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig.
-4.
-Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second.
-Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background.
-These Bozza et al.
-results buttress our case that the SN 1987A neutrinos were mostly emitted during a 0.5 to 1.0 sec time window, and that our use of $\pm 0.5$ sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified.
-A 0.5 to 1.0 sec time window is apparently small enough not to badly degrade the fit of 37 neutrino data points to three straight lines.
-However, the use of as many as 9 free parameters in the Bozza spectral fit is a serious drawback, since with enough free parameters one can fit virtually any shape.
-In contrast, our fit to the $m^{2}>0$ neutrino spectrum used only two parameters, i.e., the two $m^{2}>0$ masses.
-Moreover, as long as the spread in neutrino emission times is small compared to the horizontal error bars used (as implied by Fig.
-2), relativistic kinematics requires that neutrinos of a given mass lie on straight lines having a specific slope, and there is no similar reqirement in the Bozza analysis.
 Figure 4: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al.
-(Bozza2025, ).
+(Bozza2025, ), who use a two phase model.
+Bozza’s accretion phase luminosity is in red, and their cooling phase is in blue.
 Note the time scale change at $t=1$ seconds.
-The fraction of the integrated luminosity for $t<1$ seconds is only a bit more than half because of that scale change.
 Figure is used under creative commons license https://creativecommons.org/licenses/by/4.0
@@
 The current upper limit on the total sum of neutrino masses from cosmology has tightened to $\sum m_{\nu}<0.07-0.12eV/c^{2}$ (Shao2025, ), but some analyses of the data even favor a negative value for the sum of the neutrino masses.
-(Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found in the SN 1987A data, which could be avoided provided the mass sum were over flavor state masses not mass states.
+(Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found from the SN 1987A data, which could be avoided provided the mass sum were over flavor state masses not mass states.
 In the recent literature the issue of whether the sum is over mass or flavor states is not widely discussed.
@@
 IV.10 A “missing” $m_{4}=2.70eV/c^{2}$ sterile neutrino?
-KATRIN recently reported finding no evidence for sterile neutrinos in its data based on the observed shape of the tritium beta decay spectrum.
+In 2025 KATRIN reported finding no evidence for sterile neutrinos based on the observed shape of the tritium beta decay spectrum.
 (KATRIN2025, ).
@@
 Moreover, the conflict of KATRIN with the Neutrino-4 experiment was not a small one, since the data of the GALLEX, SAGE, and BEST experiments confirms the parameters of neutrino oscillations $\Delta m^{2}_{14}=7.3eV^{2}$ (or $m_{4}=2.70eV/c^{2}$ and $\sin^{2}2\theta_{14}=0.36)$ and the confidence level of the Neutrino-4 sterile neutrinos combined with the data of the GALLEX, SAGE, and BEST experiments is now $5.8\sigma.$ (Serebrov2024, )
-IV.11 “The dog that didn’t bark” (the key section!)
-It is possible that KATRIN failed to see any sterile neutrinos because they simply do not exist, but here we consider a much more interesting possibility.
+IV.11 Why KATRIN failed to see sterile neutrinos?
+It is possible that KATRIN failed to detect any sterile neutrinos because they do not exist, but here we consider a much more interesting possibility.
 Neutrino - 4 and the other three experiments that saw the $2.70eV/c^{2}$ neutrino were doing a short baseline disappearance oscillation experiment, which observed the disappearance probability $P(\bar{\nu_{e}}\rightarrow\bar{\nu_{e}}).$ This type of experiment would work just as well if the disappearing neutrinos were disappearing into sterile or active neutrinos.